### v1

# Quantum geometry from the Moyal product: quantum kinetic equation and non-linear response

Abstract
We systematically derive the dissipationless quantum kinetic equation for a multi-band free fermionic system with U(1) symmetry. Using the Moyal product formalism, we fully band-diagonalize the dynamics. Expanding to the second order in gradients, which is beyond the semiclassical limit, we give a complete analysis of the band-resolved thermodynamics and transport properties, especially those arising from the quantum geometric tensor. We apply our framework to a Bloch band theory under electric fields near equilibrium and find the linear and nonlinear transport coefficients. We also obtain the dynamical density-density response functions in the metallic case, including quantum metric corrections. Our results and approach can be applied very generally to multi-band problems even in situations with spatially varying Hamiltonians and distributions.
1 Introduction
Kinetic equations describe the time evolution of the phase space density of an ensemble of particles. Such equations have quantum and classical versions, and can provide a means to investigate transport and responses of many body systems. Kinetic equations provide an alternative to the method of Kubo formulae, which are formally exact perturbative calculations of time-dependent response to applied forces. Compared to the latter, kinetic equations have some advantages: they are more intuitive, can more easily treat inhomogeneous systems, and directly yield the non-equilibrium distribution function of the system under study.
While the most famous kinetic equation is the semiclassical Boltzmann equation for phase space densities, capturing the quantum coherence of particles requires using and “diagonalizing” quantum kinetic equations which involve full density matrices. Although these equations can be formulated exactly in phase phase using the Moyal formalism [1], diagonalizing them exactly is usually not possible. However, this formalism is well suited to the case of small gradients, for which one can systematically expand the equations order by order in gradients, subsequently diagonalize them, and express observables using “single-band quantities.” Under this procedure, the quantum coherence effects, in particular the multiple-band effects, manifest in each individual band through “geometric terms”, which involve derivatives of the diagonalization matrix (and in turn eigenstates of all bands). At first order in the gradients, one recovers the usual Boltzmann’s equation with the proper velocity and force using this formalism [2, 3, 4, 5]. At that order, the real, momentum and mixed space components of the Berry curvature enter. Identifying some coefficients of phase space derivatives of the distribution function with phase space velocities, one can also recover the single particle equations of motion of the center of a wavepacket in the “wavepacket formalism” [6], which can be further formulated as symplectic mechanics [7, 8, 9, 10].
In addition to the Berry curvature, the quantum metric describes another fundamental aspect of quantum geometry. It arises as the real part of the quantum geometric tensor (while the Berry curvature constitutes its imaginary part) [11, 12, 13, 14]. The quantum metric quantifies the distance between neighboring quantum states and plays a crucial role in non-adiabatic responses. Although identifying physical observables that depend solely on the quantum metric remains challenging, significant efforts have been made to characterize quantum-metric-induced linear and nonlinear responses, employing both perturbation theory and the wavepacket formalisms [15, 16, 17, 18].
Here, we push the Moyal product formalism to second order in the gradients for particle-conserving systems of non-interacting fermions without assuming translation symmetry. We show that the Berry curvature gains corrections at this order, and that the quantum metric tensor and a symmetric interband coherence tensor modify the kinetic equation and the observables. Our results are general and exact to second order in the gradients, and provide a unified framework that can be applied to many different problems including transport and dynamical response. Notably, using this systematic approach, we find in particular that the wavepacket formalism does not capture all contributions to response functions, and working up to second order also allows us to highlight, even at first order, an ambiguity in the definition of band diagonal quantities.
1.1 Main results in the case of a separable Hamiltonian
The general setting of this paper is a $d+1$-dimensional non-interacting fermion system with the particle-conserving Hamiltonian [^1] We define $\int_{x}\equiv\int\textrm{d}^{d}x$ and $\int_{p}\equiv\int\frac{\textrm{d}^{d}p}{(2\pi\hbar)^{d}}$ for integrals.
$\hat{\rm H}(t)=\sum_{n,m=1}^{N}\int_{x,x^{\prime}}\hat{\psi}_{n}^{\dagger}(x){\sf H}_{nm}(x,x^{\prime},t)\hat{\psi}_{m}(x^{\prime}),$ (1)
where $x=(x_{1},\ldots,x_{d})$ and $\hat{\psi}^{\dagger}_{n},\hat{\psi}_{n}$ $(n=1,\ldots,N)$ are fermionic fields that satisfy the anti-commutation relations $\left\{\hat{\psi}_{n}(x)\,,\hat{\psi}_{m}^{\dagger}(x^{\prime})\right\}=\delta_{nm}\delta^{d}(x-x^{\prime})$.
Within our formalism, this problem defined with quantum mechanical operators is mapped to a problem defined on phase space. We define a diagonalized Hamiltonian $h(x,p,t)$ and fermion distribution function $f(x,p,t)$ which are both $N\times N$ diagonal matrices defined on phase space in such a way that they satisfy the equations
${\rm N}(t)=\int_{x,p}\mathrm{tr}[f(x,p,t)],$ (2)
${\rm E}(t)=\langle{\rm\hat{H}}(t)\rangle=\int_{x,p}\mathrm{tr}[h(x,p,t)f(x,p,t)],$ (3)
where ${\rm N}$ and ${\rm E}$ are the total number of fermions and the expectation value of the Hamiltonian. Band structures, in the conventional sense, can only be defined for systems with translation symmetry, but by defining $f(x,p,t)$ and $h(x,p,t)$ in this way, we can think of the diagonal elements of each as the phase-space distribution functions and energies of a “band” even for systems without translation symmetry.
While this formalism that we will fully develop in section 2 applies to any particle-conserving non-interacting Hamiltonian of fermions, we are particularly interested in the case of a separable Hamiltonian of the form
$\hat{\rm H}(t)=\int_{x}\hat{\psi}^{\dagger}(x)\left({\sf H}_{0}(-i\hbar\nabla_{x})-eV(x,t)I_{N}\right)\hat{\psi}(x)$ (4)
that describes a translationally invariant system with Hamiltonian ${\sf H}_{0}$ probed by an external electric potential $V$. We assume ${\sf H}_{0}$ is a $N\times N$ matrix Hamiltonian with matrix eigenvalues $\{\varepsilon_{n}\}$ and eigenstates $\{|u_{n}\rangle\}$ for $n=1,\ldots N$, and we show that the diagonal distribution function $f(x,p=\hbar k,t)$ satisfies the kinetic equation
$\partial_{t}f=$ $-v\cdot\nabla_{x}f-\frac{e}{\hbar}\nabla_{x}V\cdot\nabla_{p}f-\frac{e}{\hbar}\partial_{x_{i}}f{\sf\Omega}_{ij}\partial_{x_{j}}V$
$+\frac{1}{24\hbar}\partial^{3}_{x_{i}x_{j}x_{l}}f\partial^{3}_{k_{i}k_{j}k_{l}}\varepsilon+\frac{e}{24\hbar}\partial^{3}_{k_{i}k_{j}k_{l}}f\partial^{3}_{x_{i}x_{j}x_{l}}V$
$-\partial_{x_{i}}\bigg{[}f\bigg{(}e\,{\sf t}_{ij}(\partial_{t}+v\cdot\nabla_{x})\partial_{x_{j}}V$
$\hskip 40.0pt+\frac{e^{2}}{2\hbar}\left(2\partial_{k_{j}}{\sf t}_{il}-\partial_{k_{i}}{\sf t}_{jl}\right)\partial_{x_{j}}V\partial_{x_{l}}V\bigg{)}\bigg{]}$
$+\frac{e}{2\hbar}\partial_{x_{i}}(f\partial^{2}_{x_{j}x_{l}}V)\partial_{k_{i}}{\sf g}_{jl}+\frac{e}{2\hbar}\partial^{2}_{x_{i}x_{l}}(f\partial_{x_{j}}V)\partial_{k_{j}}{\sf g}_{il}$
$-\frac{e}{\hbar}\partial^{2}_{k_{i}x_{j}}(f{\sf g}_{jl}\partial^{2}_{x_{i}x_{l}}V)+\mathcal{O}(\partial_{x}^{4}),$ (5)
where $i,j,l=1,..,d$. Let us describe (5). All quantities involved are either diagonal matrices or scalars. Hence, the equation above actually describes $N$ separate kinetic equations for each band. We will routinely use this approach to express equations for each band as a single equation using diagonal matrices. In addition, we assumed the Einstein summation convention which we will continue to adopt for the rest of the paper. It should also be noted that the kinetic equation is expressed using wave numbers $k=p/\hbar$ rather than momenta to adhere to the standard convention when working with Bloch states. In eq. 5, $v_{i}=\hbar^{-1}\partial_{k_{i}}\varepsilon$ is the group velocity; ${\sf\Omega}_{ij}={\sf\Omega}_{k_{i}k_{j}}$ is the Berry curvature; ${\sf g}_{ij}={\sf g}_{k_{i}k_{j}}$ is the quantum metric; and ${\sf t}_{ij}$ is an interband coherence term with $n$-th diagonal element
${\sf t}_{n,ij}\equiv-\sum_{m\neq n}\frac{\langle\partial_{k_{i}}u_{n}|u_{m}\rangle\langle u_{m}|\partial_{k_{j}}u_{n}\rangle}{\varepsilon_{n}-\varepsilon_{m}}+(i\leftrightarrow j).$ (6)
Existing literature refers to this quantity as the Berry connection polarizability [15, 19] or the band-normalized quantum metric [20, 18]. The first line in the kinetic equation is the usual Boltzmann equation with the Berry curvature contribution from the anomalous velocity [6]. The second line is a band-geometry independent correction to the Boltzmann equation coming from expanding to second order in gradients. The third and fourth lines describe interband effects and the last two lines describe the effects of the quantum metric.
Using a relaxation time approximation to approximate the collision integral $\mathcal{I}[f]$, we solve the kinetic equation perturbatively in $e$ to calculate the electric current; see (122) and (123) for the whole expression. At the order of linear response, we find the latter contains the usual Drude and Hall term, and additionally, a contribution coming from the quantum-metric dipole (QMD) tensor, $\partial_{k_{l}}{\sf g}_{ij}$, defined as:
$\widehat{J}_{i}^{\textrm{QMD}}(q,\omega)=e^{2}\left(\sum_{n=1}^{N}\int_{k}n_{F,n}\partial_{k_{l}}{\sf g}_{ij}\right)q_{j}q_{l}\widehat{V}(q,\omega)$ (7)
where $n_{F,n}\equiv n_{F}(\varepsilon_{n})$ is the Fermi-Dirac distribution function for the $n$-th band and $\widehat{V}$ is the Fourier transformed potential $V$ [17]. In addition, going to second-order response, we find the nonlinear Hall effect as expected [21] as well as contributions from the “band-normalized quantum metric” [18].
A similar calculation by Lapa and Hughes in Ref. [16] identified a contribution to the linear electrical response from the QMD tensor as well. Lapa and Hughes generalized the semiclassical wavepacket formalism to include effects of the quantum metric into the equations of motion of the wave packet, which they used as input to the Boltzmann equation in solving for the electric current. However, we observe that the overall prefactor as well as the permutation of indices of the QMD tensor in their result are different from eq. 7. At the same time, our results of second-order response are distinct at various terms from Ref. [18], which is also a generalization of the wavepacket formalism by quantum geometry. It is important to note that the above differences are not subject to the ambiguity in defining the current operator by divergence-free terms. We believe that this discrepancy can be explained by the fact that, in Ref. [16, 18], the Boltzmann equation was not properly generalized along with the wave packet equations of motions. In general, the semiclassical wavepacket approach is difficult to systematically generalize in contrast to the formalism we develop in this paper.
Moreover, by expanding the distribution function perturbatively in external fields, we obtain the dynamical $n$-point density correlation functions in (127a) and (127b) with both intraband and interband contributions. By integrating over the frequency, we obtain the static structure factor with (i) a term $\propto q$ as in usual single-band Fermi liquid and (ii) a term $\propto q^{3}$ due to the quantum metric in (128).
1.2 Outline of the manuscript
The main body of this paper is divided into three parts and is organized as follows. In section 2, we begin by first introducing the Wigner transformation and the Moyal product. The results in that section are obtained by expanding the Moyal product to second-order in $\hbar$. We then define the matrix-valued phase-space distribution function ${\sf F}$ (whose Moyal diagonalization will lead to $f$ introduced above) and its kinetic equation which together form the starting point of our formalism. We introduce the “Moyal diagonalization” of the Hamiltonian which gives us a definition for bands in phase-space (see however the discussion in section 2.4) and also leads to an emergent gauge structure. We show that for a time-independent Hamiltonian or for a slowly-varying time-dependent Hamiltonian, “Moyal diagonalization” transforms the kinetic equation for the matrix-valued ${\sf F}$ into a set of $N$ decoupled scalar kinetic equations where $N$ is the number of bands in phase-space. We also develop all of the necessary machinery to perform practical calculations using our formalism and derive an expression for the electric current. We make several assumptions to develop our formalism further, namely assume that the Hamiltonian is 1) non-interacting; 2) slowly varying in space which is necessary for the validity of the $\hbar$ expansion; 3) slowly varying in time which allows us to ignore interband transitions in time; 4) $U(1)$ charge symmetric; and 5) non-degenerate. As we will discuss in section 4, some of these assumptions can be dropped by generalizing our formalism.
In section 3, we apply the developed formalism to the case of a translationally-invariant system that is perturbed by an external electric potential (which we introduced earlier in section 1.1). Assuming a time-independent electric potential, we calculate the equilibrium electric current. We also solve the kinetic equation for the case of a time-dependent electric potential using the relaxation time approximation and we calculate the non-equilibrium electric current from which we can identify the contribution from the QMD. We also consider the collisionless limit in order to calculate the corrections to the two-point and three-point density correlation functions of a Fermi liquid coming from band geometry.
We conclude in section 4 with an extended discussion of different aspects of our formalism, a summary of results, and potential future applications and extensions.
2 Formalism
2.1 Setup
We begin by considering a non-interacting fermion system with the Hamiltonian from eq. 1. We define the density matrix
$\begin{split}{\sf F}_{nm}(x,x^{\prime},t)&\equiv\langle\hat{\psi}_{m}^{\dagger}(x^{\prime},t)\hat{\psi}_{n}^{\vphantom{\dagger}}(x,t)\rangle,\end{split}$ (8)
where $\hat{\psi}_{n}(x,t),\hat{\psi}_{m}^{\dagger}(x^{\prime},t)$ are defined in the Heisenberg picture. The time-dependence of the density matrix is dictated by the Liouville-von Neumann equation
$i\hbar\partial_{t}{\sf F}(x,x^{\prime},t)$ (9)
$\qquad=\int_{y}\left({\sf H}(x,y,t){\sf F}(y,x^{\prime},t)-{\sf F}(x,y,t){\sf H}(y,x^{\prime},t)\right),$
where matrix multiplication is implicit. In principle, we can solve for the density matrix using eq. 9 and calculate a variety of observables such as total energy, number (energy) density, number (energy) current, etc. with the solution, but this is an intractable task for a general matrix Hamiltonian.
In order to make the problem tractable, following [5], we use the Wigner transformation to express the density matrix and other bilocal fields such as ${\sf H}$ as functions defined over phase space. The Moyal or “star” product which appears in place of operator products is denoted by $\star$ and is defined as
=exp(i2(
x
p
x
p
=exp(i2ωαβ
α
β
[…]
where $\omega$ is an antisymmetric tensor with indices that run over both position and momentum coordinates and is defined as $\omega^{x_{i}p_{j}}=-\omega^{p_{i}x_{j}}=\delta_{ij}$ and $\omega^{x_{i}x_{j}}=\omega^{p_{i}p_{j}}=0$ [note that here we do not carry out a Wigner transformation in time but rather consider a single time]. In general, Greek letters, e.g. $\alpha,\mu=1,..,2d$, will be used to denote indices that run over phase-space coordinates and Roman letters, e.g. $i=1,..,d$, will be used to denote the indices of either position or momentum coordinates. Hereafter, we assume all functions are defined over phase space unless otherwise stated.
Typically, Moyal products are not evaluated exactly. Rather, they are evaluated by treating $\hbar$ as a small parameter and expanding the exponential as a series. Since each factor of $\hbar$ comes with a single position and momentum derivative, the power of $\hbar$ equals the number of additional position and momentum derivatives introduced in an expression after expanding the Moyal product. If the system of interest varies sufficiently slowly in space, this expansion in $\hbar$ becomes controlled. In this work, we will carry out this expansion to second order in $\hbar$. In section 3, we will choose to express our formulas using wave numbers $k=p/\hbar$ rather than momenta which leads to the loss of the $\hbar$ in the Moyal product. In that case, the expansion can be controlled by directly counting the number of position derivatives.
Under the Wigner transformation, eq. 9 becomes
$\begin{split}i\hbar\partial_{t}{\sf F}=&\left[{\sf H}\,\overset{\star}{,}\,{\sf F}\right]\end{split}$ (11)
where ${\sf F}$ and ${\sf H}$ are now matrix-valued functions defined on phase space and $\left[\cdot\,\overset{\star}{,}\,\cdot\right]$ is defined as the commutator with respect to the matrix and Moyal products. The density matrix is now defined on phase space, and we refer to ${\sf F}$ as the distribution function and the Liouville-von Neumann equation as its kinetic equation. In this work, we operate under the assumption that ${\sf F}\rightarrow 0$ as $\lvert{x}\rvert,\lvert{p}\rvert\rightarrow\infty$ so that the boundary terms of phase-space integrals vanish when integrating by parts.
Since ${\sf F}$ is matrix valued, by expanding the Moyal product in eq. 11, this kinetic equation becomes a set of coupled differential equations. We will now discuss the diagonalization of the Hamiltonian such that this kinetic equation can be transformed into a set of $N$ decoupled differential equations for $N$ scalar-valued distribution function.
2.2 Moyal diagonalization
Following Refs. [2, 4, 5], we introduce a unitary matrix $U$, which in the phase-space formalism satisfies the equation
$U\star U^{\dagger}=U^{\dagger}\star U=I_{N},$ (12)
where $I_{N}$ is the $N\times N$ identity matrix. We then define the “Moyal-diagonalization” of ${\sf H}$ as
$U^{\dagger}\star{\sf H}\star U\equiv\tilde{h}=\begin{pmatrix}\tilde{h}_{1}&&\\ &\ddots&\\ &&\tilde{h}_{N}\end{pmatrix}$ (13)
where $\tilde{h}=\tilde{h}(x,p)$ is a diagonal matrix that is a priori not a constant but a function on phase space. [^2] In general, lowercase letters will be used to denote diagonal quantities except for the Berry connection $A$ which will be introduced shortly. In this work, $\{\tilde{h}_{1},\cdots,\tilde{h}_{N}\}$ are assumed to be non-degenerate. If the Hamiltonian is translation-invariant, $\tilde{h}=\tilde{h}(p)$ is the dispersion relations of the band structure in the system. For an inhomogeneous system, $\tilde{h}=\tilde{h}(x,p)$ is a function on phase space, and it gives us an operational definition for a band structure even in the absence of translation symmetry. These bands defined on phase space can be thought of as semi-classically corrected bands. Further discussion on this can be found in section 4.2.
Just like $\tilde{h}$, in general, quantities that are associated with a single band such as the Berry curvature $\Omega_{n,\alpha\beta}$ or the quantum metric $g_{n,\alpha\beta}$, which we will introduce shortly, will usually be expressed using diagonal matrices. Hence, a trace over these diagonal quantities is a summation over all the bands.
The diagonalization equation eq. 47 can be solved by expanding the Moyal product order-by-order in $\hbar$. At zeroth order, it is
$U^{(0)\dagger}{\sf H}U^{(0)}=\tilde{h}^{(0)},$ (14)
where $U^{(0)},U^{(0)\dagger},\tilde{h}^{(0)}$ denote the solutions for the zeroth-order equation. Therefore, at lowest order, $\tilde{h}_{n}^{(0)}$ are simply the matrix eigenvalues of ${\sf H}$. By going to finite orders in $\hbar$, we can calculate corrections to these eigenvalues that come from the spatial variation of ${\sf H}$. Closed form expressions for these corrections will be derived in section 2.7.
We can take the star-unitary operators that are used to diagonalize ${\sf H}$ to define $\tilde{F}\equiv U^{\dagger}\star{\sf F}\star U$. Although ${\sf H}$ is diagonalized by $U$, $\tilde{F}$ is generally not a diagonal matrix. However only the diagonal components of $\tilde{F}$ contribute to observables such as the total particle number and energy. For convenience, we separate the diagonal and off-diagonal components of $\tilde{F}$ labeling them $\tilde{f}$ and $\tilde{\mathcal{F}}$ respectively such that $\tilde{F}=\tilde{f}+\tilde{\mathcal{F}}$.
2.3 Gauge symmetry
In diagonalizing the Hamiltonian in (13), we introduced a U(1)⊗N gauge redundancy meaning there are multiple $\star$-unitary transformations that can diagonalize the Hamiltonian. For simplicity, if we assume the diagonalized Hamiltonian does not have any degeneracies, i.e. $\tilde{h}_{n}\neq\tilde{h}_{m}$ if $n\neq m$, we define the gauge transformation as
$\begin{split}U&\to U\star\operatorname{E}^{i\theta},\\ U^{\dagger}&\to\operatorname{E}^{-i\theta}\star U^{\dagger},\end{split}$ (15)
where $\theta=\theta(x,p)$ is a diagonal matrix and $\operatorname{E}^{(\cdot)}$ is the “star-exponential” defined as
$\begin{split}\mathrm{E}^{\lambda}\equiv&1+\lambda+\frac{1}{2}\lambda\star\lambda+\cdots+\frac{1}{n!}\lambda^{\star n}+\cdots\\ =&\mathrm{e}^{\lambda}+\mathcal{O}(\hbar^{2}).\end{split}$ (16)
To first order in $\hbar$, the “star-exponential” is equivalent to the exponential function defined using regular multiplication. Under this gauge transformation, the diagonalized Hamiltonian is still diagonal, $\tilde{h}=U^{\dagger}\star{\sf H}\star U\rightarrow\operatorname{E}^{-i\theta}\star\tilde{h}\star\operatorname{E}^{i\theta}$, but $\tilde{h}$ is not gauge invariant. By expanding in powers of $\hbar$, we can show that under this gauge transformation, the diagonalized Hamiltonian $\tilde{h}$ as well as the diagonal component of the distribution function $\tilde{f}$ transform as
$\tilde{d}\rightarrow$ $\operatorname{E}^{-i\theta}\star\tilde{d}\star\operatorname{E}^{i\theta}$
$=\tilde{d}-\hbar\omega^{\alpha\beta}\partial_{\alpha}\tilde{d}\left(\partial_{\beta}\theta+\frac{\hbar}{2}\omega^{\sigma\lambda}\partial^{2}_{\beta\lambda}\theta\partial_{\sigma}\theta\right)$
$\quad+\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{d}\partial_{\beta}\theta\partial_{\lambda}\theta+\mathcal{O}(\hbar^{3}),$ (17)
where $\tilde{d}=\tilde{h}$ or $\tilde{f}$. Note that at zeroth order in $\hbar$, $\tilde{d}$ is invariant under this transformation. In (17), we only considered the diagonal components of the distribution function but the off-diagonal components are also gauge dependent. In this work, we use a tilde to clearly denote quantities that are gauge-dependent.
Since we have a gauge symmetry, the gauge field associated with this gauge symmetry is
$\Lambda_{\alpha}\equiv-iU^{\dagger}\star\partial_{\alpha}U,$ (18)
and we define the diagonal components of this gauge field as the Berry connection
$A_{\alpha}=\operatorname{diag}[\Lambda_{\alpha}].$ (19)
Under the gauge transformation eq. 15, the gauge field transforms as
$\Lambda_{\alpha}$ $\rightarrow\operatorname{E}^{-i\theta}\star\left(\Lambda_{\alpha}+i\operatorname{E}^{i\theta}\star\partial_{\alpha}\operatorname{E}^{-i\theta}\right)\star\operatorname{E}^{i\theta}$
$=e^{-i\theta}\left(\Lambda_{\alpha}+\frac{\hbar}{2}\omega^{\sigma\lambda}\left(i\partial_{\sigma}\theta\Lambda_{\alpha}\partial_{\lambda}\theta-\left\{\partial_{\sigma}\Lambda_{\alpha}\,,\partial_{\lambda}\theta\right\}\right)\right)e^{i\theta}$
$\qquad+\partial_{\alpha}\theta+\frac{\hbar}{2}\omega^{\sigma\lambda}\partial_{\sigma}\theta\partial^{2}_{\alpha\lambda}\theta+\mathcal{O}(\hbar^{2}),$ (20)
and the Berry connection transforms as
$A_{\alpha}$ $\rightarrow\operatorname{E}^{-i\theta}\star\left(A_{\alpha}+i\operatorname{E}^{i\theta}\star\partial_{\alpha}\operatorname{E}^{-i\theta}\right)\star\operatorname{E}^{i\theta}$
$=A_{\alpha}+\partial_{\alpha}\theta-\hbar\omega^{\sigma\lambda}\partial_{\sigma}A_{\alpha}\partial_{\lambda}\theta$
$\qquad+\frac{\hbar}{2}\omega^{\sigma\lambda}\partial_{\sigma}\theta\partial^{2}_{\alpha\lambda}\theta+\mathcal{O}(\hbar^{2}).$ (21)
Now that we have an understanding of the gauge structure, let us construct some basic gauge-invariant quantities that will come in handy later. Firstly, the simplest gauge-invariant object we can consider is the band projection operator. We define the semiclassical-band projection operator $P_{n}$ of the $n$-th semiclassical band as
$P_{n}\equiv U\star p_{n}\star U^{\dagger},$ (22)
where $p_{n}$ is a diagonal matrix with matrix elements $(p_{n})_{ij}=\delta_{ni}\delta_{nj}$ [22]. One can check explicitly that $P_{n}$ is invariant under the gauge transformation eq. 15. It also follows from the unitarity of $U$ that the projection operator is star-idempotent and complete: $P_{n}\star P_{m}=\delta_{nm}P_{n}$, $\sum_{n=1}^{N}P_{n}=I_{N}$.
In analogy with the quantum geometric tensor (QGT) defined in band theory [11, 12, 13, 14], we define the phase-space QGT as
$\begin{split}\mathcal{T}_{n,\alpha\beta}\equiv&\mathrm{tr}\left[\partial_{\alpha}P_{n}\star(I-P_{n})\star\partial_{\beta}P_{n}\right]\\ =&g_{n,\alpha\beta}+\frac{i}{2}\Omega_{n,\alpha\beta}.\end{split}$ (23)
The symmetric part which is real is the phase-space quantum metric $g_{n,\alpha\beta}$, and the antisymmetric part which is imaginary is the phase-space Berry curvature $\Omega_{n,\alpha\beta}$. For brevity, we simply refer to them as the quantum metric and Berry curvature.
From the definition above, the Berry curvature and quantum metric must be
$\begin{split}\Omega_{\alpha\beta}=&\partial_{\alpha}A_{\beta}-\partial_{\beta}A_{\alpha}-\hbar\omega^{\sigma\lambda}\partial_{\sigma}A_{\alpha}\partial_{\lambda}A_{\beta}\\ &+\hbar\omega^{\sigma\lambda}\partial_{\sigma}((\partial_{\alpha}A_{\beta}-\partial_{\beta}A_{\alpha})A_{\lambda})+\mathcal{O}(\hbar^{2}),\end{split}$ (24)
and
$g_{\alpha\beta}=\frac{1}{2}\operatorname{diag}(\Lambda_{\alpha}\Lambda_{\beta}+\Lambda_{\beta}\Lambda_{\alpha})-A_{\alpha}A_{\beta}+\mathcal{O}(\hbar).$ (25)
We only derived the quantum metric to zeroth order in $\hbar$ since it only appears at that order in observables and the kinetic equation at second order in $\hbar$ as we will see later. In addition, notice that to calculate the Berry curvature $\Omega$ at first in $\hbar$, in eq. 24 we must use the the Berry connection $A$ up to first-order. In section 2.7, we provide a formula, eq. 43, to calculate this term.
2.4 Gauge-invariant $h,f$
As we discussed in the previous section, the diagonal quantities $\tilde{h}$ and $\tilde{f}$ are gauge dependent. However, we can define a gauge-invariant form of these quantities which we denote $h,f$ respectively. We make the following choice for the gauge-invariant distribution function $f$
$\begin{split}f\equiv&\tilde{f}+\hbar\omega^{\alpha\beta}\partial_{\alpha}(\tilde{f}A_{\beta})\\ &+\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}(\tilde{f}\operatorname{diag}(\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\}))+\mathcal{O}(\hbar^{3}),\end{split}$ (26)
and we choose the gauge-invariant diagonalized Hamiltonian $h$ to be
$h\equiv$ $\tilde{h}+\hbar\omega^{\alpha\beta}\partial_{\alpha}\tilde{h}(A_{\beta}+\hbar\omega^{\sigma\lambda}A_{\lambda}\partial_{\sigma}A_{\beta})$
$-\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{h}\left(\operatorname{diag}(\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\})-4A_{\beta}A_{\lambda}\right)+\mathcal{O}(\hbar^{3}).$ (27)
Although it is tedious, eqs. 17, LABEL:eq:Lambdatransform and 21 can be used to show that $f,h$ are gauge-invariant.
As mentioned in the introduction eqs. 2 and 3 the gauge-invariant distribution function $f$ is defined such that the total number of fermions is
${\rm N}=\int_{x,p}\mathrm{tr}[{\sf F}]=\int_{x,p}\mathrm{tr}[f],$ (28)
and $h$ is defined such that the total energy is given by
${\rm E}=\langle\hat{H}\rangle=\int_{x,p}\frac{1}{2}\mathrm{tr}[\{{\sf F}\overset{\star}{,}{\sf H}\}]=\int_{x,p}\mathrm{tr}[fh].$ (29)
Note, the off-diagonal components of $\tilde{F}$, $\tilde{\mathcal{F}}$, do not contribute to either of these quantities. The details of how eqs. 26 and 27 are obtained are outlined in appendix C.
Let us finally emphasize the following point. While $f$ and $h$ defined in eqs. 26 and 27 are gauge invariant and satisfy eqs. 28 and 29, they do not constitute a unique such choice and may therefore not unequivocally represent the band filling and band energy, respectively. This is a manifestation of the fact that there exist no true “bands” when the system is not translationally invariant because momentum is in that case not a good quantum number. This notion of choice may seem surprising given the familiar semiclassical textbook equations. Indeed, up to first order in $\hbar$, some gauge-invariant choices for $f$ and $h$ are such that their relation to $\tilde{f}$ and $\tilde{h}$ can be recast in the form of a change of variables, $f(x,p)=\tilde{f}(x_{i}+\hbar A_{p_{i}},p_{i}-\hbar A_{x_{i}})$ (see e.g. Ref. [5]). In that case, ${\rm N}\sim\int{\rm tr}[\mathfrak{J}f]$, with $\mathfrak{J}=1+\hbar\Omega_{x_{i}p_{i}}$ the Jacobian of this change of variables. One may also define “single particle” quantities such as $\left[\frac{dx}{dt}\right]$ and $\left[\frac{dk}{dt}\right]$ which directly enter the kinetic equation, which takes the simple form $\partial_{t}f+\hbar\left[\frac{dx_{i}}{dt}\right]\partial_{x_{i}}f+\hbar\left[\frac{dk_{i}}{dt}\right]\partial_{k_{i}}f=0$, thereby justifying a center-of-mass and mean-momentum wavepacket analysis and the picture of a single particle evolving adiabatically in the “original” (translationally invariant) bands with some corrections. This makes this choice of $f,h$ natural, but we emphasize that even in that case it remains a choice. Note that the choices we make in this manuscript, eqs. 26 and 27 do not reduce to the form of a change of variables when truncated to first order, because there exists no such “change of variable” simplification at second order, and we instead impose the constraints eqs. 28 and 29, which allow, at least at the level of these equations, the interpretation of $f$ as a “band” density which integrates to the total number of particles and consequently $h$ a local energy density such that the integral of $fh$ is the total energy.
2.5 Kinetic equation
In this section, we discuss the kinetic equation of the distribution function. From eq. 9, $\tilde{F}=U^{\dagger}\star{\sf F}\star U$ satisfies the kinetic equation
$i\hbar\partial_{t}\tilde{F}=[\tilde{h}+\hbar\Lambda_{t}\,\overset{\star}{,}\,\tilde{F}],$ (30)
where $\Lambda_{t}\equiv-iU^{\dagger}\star\partial_{t}U$. For time-independent Hamiltonians and for a specific class of time-dependent Hamiltonians that can be expressed as ${\sf H}(t)=\sum_{n=1}^{N}P_{n}\star h_{n}(t)\star P_{n}$ where $\{P_{n}\}$ are time-independent projection operators and $\{h_{n}(t)\}$ are time-dependent scalar functions (c.f. appendix H), $\Lambda_{t}=0$. In this case, if $\tilde{F}$ is diagonal at some past time, it will remain diagonal at later times. However, here, we assume $\Lambda_{t}\neq 0$ and proceed assuming $\tilde{F}$ is generally non-diagonal. eq. 30 can be separated into two coupled kinetic equations for the diagonal ($\tilde{f}$) and off-diagonal ($\tilde{\mathcal{F}}$) components of $\tilde{F}$:
$i\hbar\partial_{t}\tilde{f}=$ $[\tilde{h}+\hbar A_{t}\,\overset{\star}{,}\,\tilde{f}]+\hbar\operatorname{diag}([\Lambda_{t}\,\overset{\star}{,}\,\tilde{\mathcal{F}}]),$ (31)
$i\hbar\partial_{t}\tilde{\mathcal{F}}=$ $[\tilde{h}+\hbar\Lambda_{t}\,\overset{\star}{,}\,\tilde{\mathcal{F}}]-\hbar\operatorname{diag}([\Lambda_{t}\,\overset{\star}{,}\,\tilde{\mathcal{F}}])$ (32)
$+\hbar[\Lambda_{t}-A_{t}\,\overset{\star}{,}\,\tilde{f}].$
From eq. 31, the kinetic equation for the gauge-invariant diagonal distribution function $f$ is
$\partial_{t}f=$ $\frac{1}{i\hbar}[h\,\overset{\star}{,}\,f]-\partial_{\alpha}\left[\omega^{\alpha\beta}f\left(\hbar(\omega^{\sigma\lambda}\partial_{\sigma}h\Omega_{\lambda\beta}-\Omega_{t\beta})\left(1-\frac{\hbar}{2}\omega^{\mu\nu}\Omega_{\mu\nu}\right)+\frac{\hbar^{2}}{2}\omega^{\sigma\lambda}\omega^{\mu\nu}\partial^{2}_{\sigma\mu}h\partial_{\beta}g_{\nu\lambda}\right)\right]$ (33)
$-\partial^{2}_{\alpha\sigma}\left[\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}f\left(\frac{1}{2}\omega^{\mu\nu}\partial_{\mu}h\partial_{\nu}g_{\beta\lambda}+\omega^{\mu\nu}\partial^{2}_{\mu\beta}hg_{\nu\lambda}-\frac{1}{2}\partial_{t}g_{\beta\lambda}\right)\right]$
$-i\operatorname{diag}\left([\Lambda_{t}\,\overset{\star}{,}\,\tilde{\mathcal{F}}]+\hbar\omega^{\alpha\beta}\partial_{\alpha}([\Lambda_{t}\,\overset{\star}{,}\,\tilde{\mathcal{F}}]A_{\beta})+\frac{\hbar^{2}}{4}\partial^{2}_{\alpha\sigma}([\Lambda_{t}\,\overset{\star}{,}\,\tilde{\mathcal{F}}]\operatorname{diag}(\left\{\Lambda_{\beta}\,\overset{\star}{,}\,\Lambda_{\lambda}\right\}))\right)+\mathcal{O}(\hbar^{3}),$
where $\Omega_{t\beta}$ is defined by simply extending our definition of the Berry curvature to also include time as an index. The last line of (33) comes from the off-diagonal components of the distribution function.
eq. 32 and eq. 33 form a complete set of kinetic equations, and in general, both must be solved simultaneously. However, in some cases, we can take a limit that that allows us to ignore the off-diagonal components of the distribution function. For example, in section 3, we will consider a translationally-invariant system that is perturbed by an external electric potential. For frequencies much smaller than the band gap nearest the Fermi energy, $\omega\ll\omega_{\textrm{gap}}$, the off-diagonal components of the distribution function are negligible, and only eq. 33 which becomes a set of $N$ decoupled differential equations needs to be solved. On the other hand, for phenomena that involve optical interband transitions such as the shift current or injection current, the off-diagonal components are crucial and cannot be ignored. In this work, we will primarily assume the off-diagonal components are negligible and leave the study of a general non-diagonal distribution function to future works.
Under this assumption, the last line in eq. 33 can be ignored. Even in that case we note that terms proportional to $f$ as well as higher-order derivatives of $f$ emerge which do not fit the form of a linear partial differential equation, i.e. $c_{a}\partial_{a}f=0$ where $a$ runs over whole spacetime. The particular form of the kinetic equation in eq. 33 is due to the choice of the gauge-invariant diagonal Hamiltonian and distribution functions defined in eqs. 27 and 26. Nevertheless, in general, the kinetic equation cannot be written in the form of a collisionless Boltzmann equation because the phase-space flow of a quantum system is compressible and does not obey Liouville’s theorem [23].
2.6 Equilibrium distribution function $f_{\textrm{eq}}$
In the previous section we derived the kinetic equation for the gauge-invariant distribution function $f$ which can be used to calculate the distribution function out of equilibrium. However, the exact, Moyal-diagonalized distribution function at equilibrium is already non-trivial at second order in $\hbar$, and is necessary for evaluating equilibrium quantities, as well as to provide a starting point for an expansion near-equilibrium. In this section, we will derive the equilibrium distribution function and show that it is also diagonal in the basis in which the Hamiltonian is diagonal.
The equilibrium distribution function in the original (“orbital”) basis at temperature $T=1/(k_{B}\beta)$ can be expressed using the imaginary time Green’s function as
${\sf F}_{\textrm{eq}}(x,p)=\frac{1}{\beta}\sum_{i\omega_{n}}e^{i\omega_{n}0^{+}}{\sf G}(x,p;i\omega_{n}),$ (34)
where $\sf G$ is the imaginary-time Green’s function in the Wigner representation and the sum is over fermionic Matsubara frequencies, $\omega_{n}=2\pi(n+1/2)/\beta$. ${\sf G}$ itself can be formally defined by the equations
$\begin{split}\left(i\omega_{n}-{\sf H}+\mu\right)\star{\sf G}(x,p;i\omega_{n})=&I_{N},\\ {\sf{G}}(x,p;i\omega_{n})\star\left(i\omega_{n}-{\sf H}+\mu\right)=&I_{N}.\end{split}$ (35)
If we now apply the star-unitary operator $U$ which is used to diagonalize the Hamiltonian and define $\tilde{G}\equiv U^{\dagger}\star{\sf G}\star U$, these equations become
$\begin{split}(i\omega_{n}-\tilde{h}+\mu)\star\tilde{G}(x,p;i\omega_{n})=&I_{N}\\ \tilde{G}(x,p;i\omega_{n})\star(i\omega_{n}-\tilde{h}+\mu)=&I_{N}.\end{split}$ (36)
Since $i\omega_{n}-\tilde{h}+\mu$ is diagonal, $\tilde{G}$ and the equilibrium distribution function must be diagonal in this basis.
It is straightforward to solve for $\tilde{G}$ by perturbatively expanding in powers of $\hbar$. The distribution function can be obtained by summing the Green’s functions over the Matsubara frequencies $\tilde{f}_{\textrm{eq}}=\beta^{-1}\sum_{i\omega_{n}}e^{i\omega_{n}0^{+}}\tilde{G}(x,p;i\omega_{n})$. This gives us
$\tilde{f}_{\textrm{eq}}=$ $n_{F}(\tilde{h})-\frac{\hbar^{2}}{16}n_{F}^{\prime\prime}(\tilde{h})\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{h}\partial^{2}_{\beta\lambda}\tilde{h}$ (37)
$-\frac{\hbar^{2}}{24}n_{F}^{\prime\prime\prime}(\tilde{h})\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{h}\partial_{\beta}\tilde{h}\partial_{\lambda}\tilde{h}+\mathcal{O}(\hbar^{3}),$
where $n_{F}(x)=1/(e^{\beta(x-\mu)}+1)$ is the Fermi-Dirac distribution function. Using eq. 26 and eq. 27, the gauge-invariant form of the equilibrium distribution function is
$f_{\textrm{eq}}=$ $\left(1+\frac{\hbar}{2}\omega^{\alpha\beta}\left(\Omega_{\alpha\beta}+\hbar\omega^{\sigma\lambda}\left(\frac{1}{2}\Omega_{\alpha\sigma}\Omega_{\beta\lambda}-\frac{1}{4}\Omega_{\alpha\beta}\Omega_{\sigma\lambda}+\partial^{2}_{\alpha\sigma}g_{\beta\lambda}\right)\right)\right)n_{F}+\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}n_{F}^{\prime}\partial_{\sigma}(\partial_{\alpha}hg_{\beta\lambda})$
$+\frac{\hbar^{2}}{2}n_{F}^{\prime\prime}\omega^{\alpha\beta}\omega^{\sigma\lambda}\left(\partial_{\alpha}h\partial_{\sigma}hg_{\beta\lambda}-\frac{1}{8}\partial^{2}_{\alpha\sigma}h\partial^{2}_{\beta\lambda}h\right)-\frac{\hbar^{2}}{24}n_{F}^{\prime\prime\prime}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}h\partial_{\beta}h\partial_{\lambda}h+\mathcal{O}(\hbar^{3}).$ (38)
where $n_{F}=n_{F}(h)$ We provide the details of the calculation leading to (38) in appendix E. $f_{\rm eq}$ in (38) solves (33) with $\partial_{t}f_{\rm eq}=0$ (equilibrium) and $\Lambda_{t}=0$.
2.7 Gauge-invariant diagonalized Hamiltonian $h$
In section 2.2, we introduced the Moyal diagonalization of the Hamiltonian defined by eq. 13. As we discussed, this equation can be solved by expanding the Moyal product in powers of $\hbar$. To zeroth order in $\hbar$, the diagonalization equation becomes a matrix diagonalization problem, and the diagonalized Hamiltonian is a diagonal matrix formed by the matrix eigenvalues of the Hamiltonian, $\tilde{h}^{(0)}_{n}$. In addition, as we mentioned in section 2.3, this zeroth order contribution is gauge-invariant, $\tilde{h}^{(0)}_{n}=h^{(0)}_{n}$. By going to finite orders of $\hbar$, we can obtain corrections to $\tilde{h}^{(0)}$ that come from the variation of the Hamiltonian in phase space. To keep track of the order of these corrections, we introduce the notation
$\begin{split}\tilde{h}&=\tilde{h}^{(0)}+\tilde{h}^{(1)}+\cdots,\\ \Lambda_{\alpha}&=\Lambda_{\alpha}^{(0)}+\Lambda_{\alpha}^{(1)}+\cdots,\end{split}$ (39)
where $\tilde{h}^{(n)},\Lambda_{\alpha}^{(n)}\sim\hbar^{n}$. As we mentioned earlier, the power of $\hbar$ counts the number of position and momentum derivatives, so $\tilde{h}^{(n)},\Lambda^{(n)}_{\alpha}$ are correction terms with $n$ additional position and momentum derivatives. This notation will also be applied to other quantities. In this section, we compute explicit expressions of the corrections to $h^{(0)}=\tilde{h}^{(0)}$.
One convenient tool to expand the diagonalization equation eq. 13 and calculate the corrections is what we call “flow equations” which we introduce in appendix D, but here, we simply present our results. For simplicity, we assume the matrix eigenvalues $\tilde{h}^{(0)}$ are non-degenerate. For the $n$-th band, we define the tensor $m_{n,\alpha\beta}$
$m_{n,\alpha\beta}\equiv\operatorname{Im}\mathrm{tr}[\partial_{\alpha}P_{n}\star({\sf H}-h_{n})\star\partial_{\beta}P_{n}].$ (40)
which defines a “full” orbital magnetization, i.e. an extension of the usual orbital magnetization to the Moyal formalism. Then, the first-order correction to the gauge-invariant diagonalized Hamiltonian written as a diagonal matrix is
$h^{(1)}=\frac{\hbar}{2}\omega^{\alpha\beta}m^{(0)}_{\alpha\beta}=\frac{i\hbar}{4}\omega^{\alpha\beta}\operatorname{diag}(\{\Lambda_{\alpha}^{(0)}\,,[\Lambda_{\beta}^{(0)}\,,h^{(0)}]\}).$ (41)
The second order correction is
$h^{(2)}=\frac{\hbar}{4}\omega^{\alpha\beta}m^{(1)}_{\alpha\beta}+\frac{\hbar^{2}}{16}\omega^{\alpha\beta}\omega^{\sigma\lambda}h^{(0)}\Omega_{\alpha\beta}^{(0)}\Omega_{\sigma\lambda}^{(0)}$
$+\frac{\hbar}{2}\omega^{\alpha\beta}\partial_{\alpha}h^{(0)}\big{(}A_{\beta}^{(1)}+\hbar\omega^{\sigma\lambda}\big{(}\partial_{\sigma}A_{\beta}^{(0)}A_{\lambda}^{(0)}+\frac{1}{2}A_{\sigma}^{(0)}\partial_{\beta}A_{\lambda}^{(0)}\big{)}\big{)}$
$-\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}(\partial_{\alpha}h^{(0)}\partial_{\sigma}g^{(0)}_{\beta\lambda}+2\partial^{2}_{\alpha\sigma}h^{(0)}g^{(0)}_{\beta\lambda}).$ (42)
To evaluate $h^{(2)}$ and $m^{(1)}_{\alpha\beta}$ we also need to know the first order correction to the Berry connection and the projection operator. These can also be derived using their flow equations. They are given by
$\begin{split}A_{\alpha}^{(1)}=&\operatorname{diag}([\Lambda_{\alpha}^{(0)}\,,Y^{(1)}])\\ &+\frac{\hbar}{4}\omega^{\sigma\lambda}\operatorname{diag}(\{\Lambda_{\sigma}^{(0)}\,,\partial_{\lambda}\Lambda_{\alpha}^{(0)}\}),\end{split}$ (43)
$P_{n}^{(1)}=$ $U^{(0)}\left([Y^{(1)}\,,p_{n}]+\frac{\hbar}{4}\omega^{\alpha\beta}\{\partial_{\alpha}\Lambda^{(0)}_{\beta}\,,p_{n}\}\right)U^{(0)\dagger}$
$+\frac{i\hbar}{2}\omega^{\alpha\beta}U^{(0)}\Lambda_{\alpha}^{(0)}p_{n}\Lambda^{(0)}_{\beta}U^{(0)\dagger},$ (44)
where $Y^{(1)}$ is a matrix with only off-diagonal elements given by
$(Y^{(1)})_{nm}=$ (45)
$\hskip 30.0pt\hbar\,\omega^{\alpha\beta}\frac{(\{\Lambda_{\beta}^{(0)}\,,\partial_{\alpha}h^{(0)}\}-\frac{i}{2}\{\Lambda^{(0)}_{\alpha}\,,[\Lambda^{(0)}_{\beta}\,,h^{(0)}]\})_{nm}}{2(h^{(0)}_{n}-h^{(0)}_{m})}.$
$A_{\alpha}^{(0)}$ is simply $A_{\alpha}^{(0)}={\rm diag}\Lambda_{\alpha}^{(0)}$.
2.8 Electric current density
In section 2.4, we discussed how the total fermion number and energy can be obtained from the gauge-invariant diagonalized Hamiltonian and distribution function $h$ and $f$. Another physical observable of interest is the electric current density. In this section, we derive a general expression for the electric current density (in real space, “momentum-integrated”) and find an explicit expression of the contributions to the current coming from only the diagonal components of the distribution function.
The electric current density can be obtained by introducing an external $U(1)$ gauge field $\mathcal{A}_{i}(x)$ and taking the derivative of the Hamiltonian with respect to the gauge field, i.e. $J_{i}(x)\equiv-\delta\langle{\rm\hat{H}}\rangle/\delta\mathcal{A}_{i}(x)$ (see appendix F). What we find to second order in $\hbar$ is
$J_{i}(x)=$ $-e\int_{p}\operatorname{Re}\mathrm{tr}\left[\frac{\partial{\sf H}(x,\pi)}{\partial p_{i}}\star{\sf F}(x,p)\right]$ (46)
$-\frac{e\hbar^{2}}{12}\frac{\partial^{2}}{\partial x_{k}\partial x_{l}}\int_{p}\mathrm{tr}\left[\frac{\partial^{3}{\sf H}(x,\pi)}{\partial p_{i}\partial p_{k}\partial p_{l}}{\sf F}\right]+\mathcal{O}(\hbar^{3}),$
where $\pi_{i}(x,p)\equiv p_{i}+e\mathcal{A}_{i}(x)$. For simplicity, we turn off the external gauge field $\mathcal{A}=0$.
Assuming the off-diagonal components $\tilde{\mathcal{F}}$ are negligible, we can express the current density using the diagonalized Hamiltonian $h$ and the diagonal components of the distribution function $f$. The details of this calculation are provided in appendix I. We find that the current density can be expressed as
$J_{i}(x)=$ $-e\int_{p}\mathrm{tr}\left[f\left(\partial_{p_{i}}h-\hbar\omega^{\alpha\beta}\partial_{\alpha}h\,\Omega_{p_{i}\beta}\left(1-\hbar\Omega_{x_{j}p_{j}}\right)+\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}h\partial_{p_{i}}g_{\beta\lambda}\right)\right]$
$-e\hbar\partial_{x_{j}}\int_{p}\mathrm{tr}[f(b_{ij}^{s}+b_{ij}^{a})]-\frac{e\hbar^{2}}{12}\partial^{2}_{x_{j}x_{k}}\int_{p}\mathrm{tr}[f(b^{s}_{ijk}+2b^{m}_{ij;k})]+\mathcal{O}(\hbar^{3}),$ (47)
where $b^{s}_{ij}$ and $b^{a}_{ij}$ are symmetric and anti-symmetric tensors defined as
$b_{ij}^{s}\equiv\frac{\hbar}{2}\omega^{\sigma\lambda}(\partial_{\sigma}h\partial_{\lambda}g_{p_{i}p_{j}}+\partial^{2}_{\sigma p_{i}}hg_{p_{j}\lambda}+\partial^{2}_{\sigma p_{j}}hg_{p_{i}\lambda}),$ (48)
$b_{ij}^{a}$ $\equiv-m_{p_{i}p_{j}}\left(1-\frac{\hbar}{2}\omega^{\sigma\lambda}\Omega_{\sigma\lambda}\right)$ (49)
$+\frac{\hbar}{2}\omega^{\sigma\lambda}(\partial_{p_{i}}(\partial_{\sigma}hg_{p_{j}\lambda})-\partial_{p_{j}}(\partial_{\sigma}hg_{p_{i}\lambda})),$
where $m_{p_{i}p_{j}}$ was defined in eq. 40. $b^{s}_{ijk}$ is a fully symmetric tensor defined as
$b_{ijk}^{s}\equiv-\frac{1}{2}\partial^{3}_{p_{i}p_{j}p_{k}}h,$ (50)
and $b^{m}_{ij;k}$ is a tensor with mixed symmetry defined as
$b_{ij;k}^{m}\equiv\partial_{p_{i}}c_{jk}-\partial_{p_{j}}c_{ik}+2(c_{jk;i}-c_{ik;j})$ (51)
and satisfies $b^{m}_{ij;k}=-b^{m}_{ji;k}$. Here, $c_{ik},c_{ik;j}$ are diagonal matrices with $n$-th diagonal elements
$c_{n,ij}$ $\equiv\operatorname{Re}\mathrm{tr}\left[\partial_{p_{i}}P_{n}\star({\sf H}-h_{n})\star\partial_{p_{j}}P_{n}\right],$ (52)
$c_{n,ik;j}$ $\equiv\operatorname{Re}\mathrm{tr}\left[\partial_{p_{i}p_{k}}^{2}P_{n}\star({\sf H}-h_{n})\star\partial_{p_{j}}P_{n}\right].$ (53)
The contributions to current density from the terms $b^{a}_{ij},b^{m}_{ij;k}$ are divergenceless, so they correspond to bound currents. $c_{ij}$, which appears in $b_{ij;k}^{m}$, has been referred to as the “band Drude weight” in existing literature [24] and its momentum integral over occupied states has been shown to be the correction to the free-electron Drude weight coming from the periodic potential of a crystal [25, 26].
By substituting the expression for the equilibrium distribution function eq. 38 into this expression, we can also obtain an expression for the equilibrium current density, $J_{i}^{\textrm{eq}}(x)$. The equilibrium current is divergenceless and can generally be expressed in the form $J_{i}^{\textrm{eq}}(x)=\partial_{x_{j}}\mathcal{M}_{ij}(x)$ where $\mathcal{M}_{ij}=-\mathcal{M}_{ji}$ is an antisymmetric tensor given by
$\mathcal{M}_{ij}=$ $-e\hbar\int_{p}\mathrm{tr}\bigg{[}n_{F}\bigg{(}\partial_{p_{i}}h\left(A_{p_{j}}+\frac{\hbar}{2}\omega^{\mu\nu}\partial_{\mu}(A_{p_{j}}A_{\nu})\right)+\frac{\hbar}{2}\omega^{\mu\nu}\partial_{\nu}h(A_{p_{j}}\partial_{p_{i}}A_{\mu}+A_{\mu}\partial_{p_{j}}A_{p_{i}}+A_{p_{i}}\partial_{\mu}A_{p_{j}})$ (54)
$\qquad-\frac{1}{2}m_{p_{i}p_{j}}+\frac{\hbar}{2}\omega^{\sigma\lambda}\partial_{p_{i}}(\partial_{\sigma}hg_{p_{j}\lambda})\bigg{)}\bigg{]}-\frac{e\hbar^{2}}{2}\omega^{\mu\nu}\int_{p}\mathrm{tr}[\partial_{\mu}(n_{F}g_{\nu p_{j}})\partial_{p_{i}}h]$
$+\frac{e\hbar^{2}}{24}\omega^{\mu\nu}\int_{p}\mathrm{tr}[n_{F}^{\prime}\partial^{2}_{p_{j}\mu}h\partial^{2}_{p_{i}\nu}h]-\frac{e\hbar^{2}}{12}\partial_{x_{k}}\int_{p}\mathrm{tr}[n_{F}b^{m}_{ij;k}]-(i\leftrightarrow j)+\mathcal{O}(\hbar^{3}),$
where $n_{F}=n_{F}(h)$. The expression derived above is true when there are no degeneracies in the energy. As a consequence, the equilibrium current does no work in the bulk: the integral of the Joule heating under an electric field $E_{i}=-e\partial_{x_{i}}V$ vanishes, $\int_{x}J^{\mathrm{eq}}_{i}E_{i}=e\int_{x}\partial_{x_{i}}J^{\mathrm{eq}}_{i}V=0$ where we used integration by parts.
2.9 Uniform magnetic field
A uniform magnetic field can be introduced in our formalism relatively simply by changing variables from the phase-space coordinates $(x,p)$ to $(x,\pi)$ where $\pi_{i}=p_{i}+e\mathcal{A}_{i}$ is kinematic momentum, where $\mathcal{A}_{i}$ is the external vector potential, as before. Under this change of variables, the Moyal product becomes [27, 4, 28]
$\star_{B}=$ $\exp\left(\frac{i\hbar}{2}({\mathchoice{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\displaystyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\displaystyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\textstyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\textstyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptstyle\vec{}\mkern 4.0mu$}\cr\kern-3.01389pt\cr$\scriptstyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptscriptstyle\vec{}\mkern 4.0mu$}\cr\kern-2.15277pt\cr$\scriptscriptstyle\nabla$\cr}}}}_{x}\cdot\vec{\nabla}_{\pi}-{\mathchoice{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\displaystyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\displaystyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\textstyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\textstyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptstyle\vec{}\mkern 4.0mu$}\cr\kern-3.01389pt\cr$\scriptstyle\nabla$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptscriptstyle\vec{}\mkern 4.0mu$}\cr\kern-2.15277pt\cr$\scriptscriptstyle\nabla$\cr}}}}_{\pi}\cdot\vec{\nabla}_{x})-\frac{ie\hbar}{2}\mathfrak{F}_{ij}{\mathchoice{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\displaystyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\displaystyle\partial$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\textstyle\vec{}\mkern 4.0mu$}\cr\kern-4.30554pt\cr$\textstyle\partial$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptstyle\vec{}\mkern 4.0mu$}\cr\kern-3.01389pt\cr$\scriptstyle\partial$\cr}}}{\vbox{\offinterlineskip\halign{#\cr\reflectbox{$\scriptscriptstyle\vec{}\mkern 4.0mu$}\cr\kern-2.15277pt\cr$\scriptscriptstyle\partial$\cr}}}}_{\pi_{i}}\vec{\partial}_{\pi_{j}}\right)$ (79)
[…]
where $\mathfrak{F}_{ij}=\partial_{x_{i}}\mathcal{A}_{i}-\partial_{x_{j}}\mathcal{A}_{j}$ is the constant magnetic field tensor. Here, we defined a new anti-symmetric tensor $\omega^{\alpha\beta}_{B}$ such that $\omega_{B}^{\pi_{i}\pi_{j}}\equiv-e\mathfrak{F}_{ij}$ and $\omega_{B}^{x_{i}\pi_{j}}=1$. Therefore, introducing a uniform magnetic field is as simple as making the replacement $p_{i}\rightarrow\pi_{i}$ and $\omega^{\alpha\beta}\rightarrow\omega^{\alpha\beta}_{B}$.
In appendix J, we apply this magnetic Moyal product to the equilibrium current, and find that the divergence of the latter is non-vanishing at a Weyl node due to the chiral anomaly effect [7, 29]. A more detailed study of the magnetic field contributions to second order in $\hbar$ is left for the future.
3 Application: Transport driven by electric fields
In this section, we apply the formalism developed in section 2 to study a translationally invariant system perturbed by an externally applied electric potential. We first find an expression for the equilibrium current assuming a static potential that is bounded from below. Then, we discuss the assumption of ignoring the off-diagonal components of the distribution function, $\tilde{\mathcal{F}}$, and solve the kinetic equation to find the current out of equilibrium. Lastly, we demonstrate the power of our formalism by calculating the two- and three-point density correlation functions of a collisionless Fermi liquid with non-trivial band geometry.
The Hamiltonian in the Wigner representation is given by
${\sf H}(x,p,t)={\sf H}_{0}(p)-eV(x,t)I_{N},$ (89)
where ${\sf H}_{0}(p)$ is a $N\times N$ Hamiltonian that is generally non diagonal and $-e<0$ is the electric charge of the fermion. As mentioned before, at zeroth order in $\hbar$, the diagonalization equation reduces to a matrix diagonalization problem, so the columns of $U^{(0)}$ are the eigenvectors of ${\sf H}_{0}(p)$. If the eigenvector and eigenvalues of ${\sf H}_{0}(p)$ are given by $\left\{|u_{n}(p)\rangle\right\}$ and $\left\{\varepsilon_{n}(p)\right\}$,
$U^{(0)}(p)=\bigg{[}|u_{1}(p)\rangle\,,~{}|u_{2}(p)\rangle\,,~{}\cdots\,,~{}|u_{N}(p)\rangle\bigg{]},$ (90)
$h^{(0)}(p)=\begin{pmatrix}\varepsilon_{1}(p)&&\\ &\ddots&\\ &&\varepsilon_{N}(p)\end{pmatrix},$ (91)
and the gauge field at zeroth order must be
$\left(\Lambda^{(0)}_{p_{i}}\right)_{nm}$ $=-i\langle u_{n}|\partial_{p_{i}}u_{m}\rangle,$ (92)
$\left(\Lambda^{(0)}_{x_{i}}\right)_{nm}$ $=0.$ (93)
From eq. 45, we find for $n\neq m$
$(Y^{(1)})_{nm}=-e\hbar\frac{(\Lambda_{p_{i}}^{(0)})_{nm}}{\varepsilon_{n}-\varepsilon_{m}}\partial_{x_{i}}V\,,$ (94)
and using eq. 43 we find that the first-order correction to $A^{(0)}$ is
$A^{(1)}_{n,p_{i}}=-e\hbar\,t_{n,ij}\partial_{x_{j}}V,$ (95)
where $t_{n,ij}$ is the symmetric tensor already defined in the introduction, (6),
$t_{n,ij}\equiv$  $\,2\operatorname{Re}\mathrm{tr}\left[\partial_{p_{i}}{\sf P}_{n}\frac{1-{\sf P}_{n}}{{\sf H}_{0}-\varepsilon_{n}}\partial_{p_{j}}{\sf P}_{n}\right]$ (96)
$=$ $\sum_{m\neq n}\frac{\langle\partial_{p_{i}}u_{n}|u_{m}\rangle\langle u_{m}|\partial_{p_{j}}u_{n}\rangle+(i\leftrightarrow j)}{\varepsilon_{m}-\varepsilon_{n}}.$
Here, ${\sf P}_{n}$ is the projection operator to zeroth order in $\hbar$ given by ${\sf P}_{n}\equiv P_{n}^{(0)}=|u_{n}\rangle\langle u_{n}|$. As mentioned in the introduction, $t_{n,ij}$ in existing literature is referred to as either the “Berry connection polarizability” [15, 19] or as the “band-normalized quantum metric” [20, 18]. We want to emphasize that in general this quantity is not a band-geometric quantity since it is a multi-band quantity as its calculation relies on the knowledge of all bands. Only in the case of a two-band model can it truly be considered a band-geometric quantity since the energy denominator in the summation is simply replaced by the energy difference of the two bands.
Using eq. 44 the first-order correction to this projection operator is
$P_{n}^{(1)}=-ie\hbar\partial_{x_{l}}V\left[\partial_{p_{l}}{\sf P}_{n}\,,\frac{1-{\sf P}_{n}}{{\sf H}_{0}-\varepsilon}\right].$ (97)
Next, using eq. 41 and eq. 42 we can calculate the first- and second-order corrections to $h^{(0)}$ to get
$h_{n}(x,p)=$ $\varepsilon_{n}(p)-eV(x)$ (98)
$+\frac{e^{2}\hbar^{2}}{2}t_{n,ij}(p)\partial_{x_{i}}V(x)\partial_{x_{j}}V(x)+\mathcal{O}(\hbar^{3}).$
The quantum metric is only nonzero for momentum indices and we find that
$g_{n,p_{i}p_{j}}=$ $\frac{1}{2}\left(\langle\partial_{p_{i}}u_{n}|\partial_{p_{j}}u_{n}\rangle-\langle\partial_{p_{i}}u_{n}|u_{n}\rangle\langle u_{n}|\partial_{p_{j}}u_{n}\rangle\right)$
$+\left(i\leftrightarrow j\right)+\mathcal{O}(\hbar^{2}).$ (99)
Similarly, the only finite components of the Berry curvature are
$\Omega_{p_{i}p_{j}}=$ $\partial_{p_{i}}A^{(0)}_{p_{j}}-\partial_{p_{j}}A^{(0)}_{p_{i}}$ (100)
$-e\hbar\left(\partial_{p_{i}}t_{jl}-\partial_{p_{j}}t_{il}\right)\partial_{x_{l}}V+\mathcal{O}(\hbar^{2}),$
$\Omega_{x_{i}p_{j}}=$ $-e\hbar\,t_{jl}\partial^{2}_{x_{i}x_{l}}V+\mathcal{O}(\hbar^{2}).$ (101)
The correction to the Berry curvature in eq. 100 matches the correction calculated in [18].
Assuming $V$ is time dependent, $\Lambda_{t}$ has nonzero off-diagonal components given by
$(\Lambda_{t})_{nm}=ie\hbar\frac{(\Lambda_{p_{i}}^{(0)})_{nm}}{\varepsilon_{n}-\varepsilon_{m}}\partial^{2}_{tx_{i}}V+\mathcal{O}(\hbar^{2})$ (102)
for $n\neq m$, and vanishing diagonal components.
In our work, we have taken derivatives with respect to momentum $p$ so that we can use $\hbar$ as a parameter to expand the Moyal product, but quantities such as the Berry curvature and quantum metric are typically defined using derivatives of wave number $k=p/\hbar$. Therefore, we redefine the zeroth order Berry curvature, quantum metric, and $t_{ij}$ using wave number derivatives:
${\sf\Lambda}_{i}(k)$ $\equiv\hbar\Lambda_{p_{i}}^{(0)}(\hbar k),$ (103)
${\sf A}_{n,i}(k)$ $\equiv\hbar A_{n,p_{i}}^{(0)}(\hbar k),$ (104)
${\sf\Omega}_{n,ij}(k)$ $\equiv\hbar^{2}\Omega_{n,p_{i}p_{j}}^{(0)}(\hbar k),$ (105)
${\sf g}_{n,ij}(k)$ $\equiv\hbar^{2}g_{n,p_{i}p_{j}}^{(0)}(\hbar k),$ (106)
${\sf t}_{n,ij}(k)$ $\equiv\hbar^{2}t_{n,p_{i}p_{j}}(\hbar k).$ (107)
Effectively, all we did was multiply by factors of $\hbar$ since wave number and momentum are related by $p=\hbar k$. We emphasize, that once these substitutions are made, the power of $\hbar$ no longer corresponds to the order of the Moyal product expansion. Instead, we will directly count the number of spatial derivatives.
3.1 Equilibrium current
Let us consider the case of a static external potential that is bounded from below such that an equilibrium state exists and calculate the equilibrium current. Firstly, we find that $m_{n,p_{i}p_{j}}$ defined by eq. 40 is
$m_{n,p_{i}p_{j}}=$ $\frac{1}{e\hbar}{\sf m}_{n,ij}^{\textrm{orb}}$ (108)
$-\frac{e}{\hbar^{2}}\partial_{x_{l}}V(\partial_{k_{i}}{\sf g}_{n,jl}-\partial_{k_{j}}{\sf g}_{n,il}+{\sf s}_{n,ij;l})+\mathcal{O}(\partial_{x}^{2}),$
where ${\sf m}^{\textrm{orb}}_{n,ij}$ is the anti-symmetric orbital magnetization tensor defined as [30, 31, 32]
${\sf m}^{\textrm{orb}}_{n,ij}$ $\equiv\frac{e}{\hbar}\operatorname{Im}\mathrm{tr}[\partial_{k_{i}}{\sf P}_{n}({\sf H}_{0}-\varepsilon_{n})\partial_{k_{j}}{\sf P}_{n}]$ (109)
$=-\frac{ie}{2\hbar}\langle\partial_{k_{i}}u_{n}|{\sf H}_{0}-\varepsilon_{n}|\partial_{k_{j}}u_{n}\rangle-(i\leftrightarrow j),$
and ${\sf s}_{n,ij;l}$ is defined as
${\sf s}_{n,ij;l}\equiv$ $\operatorname{Re}\mathrm{tr}\left[\left[{\sf H}_{0}-\varepsilon_{n},\partial_{k_{j}}{\sf P}_{n}\right]\partial_{k_{l}}{\sf P}_{n}\partial_{k_{i}}\left(\frac{1-{\sf P}_{n}}{{\sf H}_{0}-\varepsilon_{n}}\right)\right]$
$-(i\leftrightarrow j).$ (110)
Then, the equilibrium current is given by $J^{\textrm{eq}}_{i}(x)=\partial_{x_{j}}\mathcal{M}_{ij}$ where $\mathcal{M}_{ij}$ is (cf. (54))
$\mathcal{M}_{ij}=$ $\sum_{n=1}^{N}\int_{k}f_{n}\left(-e(v_{n,i}{\sf A}_{n,j}-v_{n,j}{\sf A}_{n,i})+{\sf m}_{n,ij}^{\textrm{orb}}\right)+\sum_{n=1}^{N}\frac{e^{2}}{\hbar}\int_{k}f_{n}\left(\hbar v_{n,i}{\sf t}_{n,jl}-\hbar v_{n,j}{\sf t}_{n,il}+\partial_{k_{i}}{\sf g}_{n,jl}-\partial_{k_{j}}{\sf g}_{n,il}+{\sf s}_{n,ij;l}\right)\partial_{x_{l}}V$
$+\sum_{n=1}^{N}\frac{e^{2}}{6\hbar}\int_{k}f_{n}^{\prime}(\partial_{k_{i}}{\sf c}_{n,jl}-\partial_{k_{j}}{\sf c}_{n,il}+2({\sf c}_{n,jl;i}-{\sf c}_{n,il;j}))\partial_{x_{l}}V+\mathcal{O}(\partial_{x}^{3}).$ (111)
Here, $f_{n}\equiv n_{F}(\varepsilon_{n}-eV)$, $v_{n,i}=\hbar^{-1}\partial_{k_{i}}\varepsilon_{n}$, and (cf. eqs. 52 and 53))
${\sf c}_{n,jl}\equiv\operatorname{Re}\mathrm{tr}[\partial_{k_{j}}{\sf P}_{n}({\sf H}_{0}-\varepsilon_{n})\partial_{k_{l}}{\sf P}_{n}],$ (112)
${\sf c}_{n,jl;i}\equiv\operatorname{Re}\mathrm{tr}[\partial^{2}_{k_{j}k_{l}}{\sf P}_{n}({\sf H}_{0}-\varepsilon_{n})\partial_{k_{i}}{\sf P}_{n}].$ (113)
The first two terms in the first line are expressed using the Berry connection which is gauge dependent. However, as mentioned before, the current itself is gauge-invariant. Indeed the derivative $\partial_{x_{j}}$ acts on $f_{n}$ to produce $-f_{n}^{\prime}e\partial_{x_{i}}V$. Then, by integration by parts, these two terms together can be rewritten as $-(e^{2}/\hbar)\sum_{n=1}^{N}\int_{k}f_{n}{\sf\Omega}_{n,ij}\partial_{x_{j}}V$ which, of course, is the Hall current. As mentioned earlier, this expression is valid when $\{\varepsilon_{n}\}$ are non-degenerate. For the case of a Weyl fermion, discussed in appendix J, the Berry connection ${\sf A}$ is ill-defined at the Weyl point, so the magnetization $\mathcal{M}_{ij}$ cannot be defined.
3.2 Non-equilibrium current
We now turn to the case of a time-dependent external potential.
3.2.1 Kinetic equation
We begin by discussing the off-diagonal components of the distribution function $\tilde{\mathcal{F}}$ which obey the kinetic equation eq. 32. We can solve for the off-diagonal components by expanding perturbatively in $e$. To first order in $e$, $\tilde{\mathcal{F}}$ must satisfy the kinetic equation
$i\hbar\partial_{t}(\tilde{\mathcal{F}})_{nm}-\varepsilon_{n}\star(\tilde{\mathcal{F}})_{nm}+(\tilde{\mathcal{F}})_{nm}\star\varepsilon_{m}$ (114)
$\approx\hbar(\Lambda_{t})_{nm}\star n_{F,m}-n_{F,n}\star\hbar(\Lambda_{t})_{nm}+\mathcal{O}(e^{2}),$
where we assumed $\tilde{\mathcal{F}}$ vanishes in the absence of an external potential. If we define the Fourier transform ~$\tilde{\mathcal{F}}(x,k,t)=\int_{q,\omega}\widehat{\tilde{\mathcal{F}}}_{nm}(q,k,\omega)e^{iq\cdot x-i\omega t}$, we find to first order in small $q$ (the expansion in $q$ is equivalent to the expansion in spatial gradients)
(~^$(\widehat{\tilde{\mathcal{F}}})_{nm}=$ $-ie\hbar\frac{({\sf\Lambda}_{i})_{nm}}{\varepsilon_{n}-\varepsilon_{m}}\frac{n_{F,n}-n_{F,m}}{\hbar\omega-\varepsilon_{n}+\varepsilon_{m}}\omega q_{i}\widehat{V}(q,\omega)$
$+\mathcal{O}(e^{2},q^{2}),$ (115)
where $n_{F,n}=n_{F}(\varepsilon_{n})$. At this order, if $\hbar\omega,eV\ll\varepsilon_{n}-\varepsilon_{m}$, we see that (~^$(\widehat{\tilde{\mathcal{F}}})_{nm}\sim\hbar\omega e\widehat{V}/(\varepsilon_{n}-\varepsilon_{m})^{2}\ll 1$. Therefore, in the low-frequency limit $\omega\ll\omega_{\textrm{gap}}$ where $\hbar\omega_{\textrm{gap}}$ is the smallest of the band gaps neighboring the Fermi energy, the contribution from $\tilde{\mathcal{F}}$ is negligible. Physically, this simply means the off-diagonal components can be ignored if we consider frequencies that are sufficiently small such that the external potential cannot induce interband transitions. This is the justification for the assumption that we made earlier that ~$\tilde{\mathcal{F}}=0$.
Now, we only have to solve the kinetic equation for $f$ given by eq. 33 which becomes (this is (5) with the addition of a collision integral $\mathcal{I}[f]$)
$\partial_{t}f=$ $-\hbar^{-1}\nabla_{k}\varepsilon\cdot\nabla_{x}f-e\hbar^{-1}\nabla_{x}V\cdot\nabla_{k}f-\frac{e}{\hbar}\partial_{x_{i}}f{\sf\Omega}_{ij}\partial_{x_{j}}V+\frac{1}{24\hbar}\partial^{3}_{x_{i}x_{j}x_{l}}f\partial^{3}_{k_{i}k_{j}k_{l}}\varepsilon+\frac{e}{24\hbar}\partial^{3}_{k_{i}k_{j}k_{l}}f\partial^{3}_{x_{i}x_{j}x_{l}}V$
$-\partial_{x_{i}}\left[f\left(e\,{\sf t}_{ij}(\partial_{t}+v_{l}\partial_{x_{l}})\partial_{x_{j}}V+\frac{e^{2}}{2\hbar}\left(2\partial_{k_{j}}{\sf t}_{il}-\partial_{k_{i}}{\sf t}_{jl}\right)\partial_{x_{j}}V\partial_{x_{l}}V\right)\right]$
$+\frac{e}{2\hbar}\partial_{x_{i}}(f\partial^{2}_{x_{j}x_{l}}V)\partial_{k_{i}}{\sf g}_{jl}+\frac{e}{2\hbar}\partial^{2}_{x_{i}x_{l}}(f\partial_{x_{j}}V)\partial_{k_{j}}{\sf g}_{il}-\frac{e}{\hbar}\partial^{2}_{k_{i}x_{j}}(f{\sf g}_{jl}\partial^{2}_{x_{i}x_{l}}V)+\mathcal{I}[f]+\mathcal{O}(\partial_{x}^{4}).$ (116)
We remind the reader that all the quantities mentioned above are diagonal matrices.
3.2.2 Definition of the relaxation time approximation
We added a collision integral to incorporate scattering effects which we will approximate using the relaxation time approximation (RTA), $\mathcal{I}[f]=-\tau^{-1}(f-f^{r})$. The RTA can be understood as modeling relaxation by coupling the system to a large external bath that maintains a distribution function $f^{r}$ and allowing fermions to scatter between the system and bath with a rate $\tau^{-1}$ (in particular, the total number of particles is not conserved). We choose the distribution function of the bath such that the system would be in thermal equilibrium with the bath in the absence of an external potential. The density matrix of the external bath in the original basis before diagonalizing is ${\sf F}^{r}=\sum_{n=1}^{N}{\sf P}_{n}n_{F,n}$. If we diagonalize this and express it in its gauge-invariant form we get the $n$-th diagonal component of $f^{r}$,
$f^{r}_{n}=(1-e\partial^{2}_{x_{i}x_{j}}V{\sf t}_{n,ij})n_{F,n}$ (117)
$-e^{2}\sum_{m\neq n}\frac{n_{F,n}-n_{F,m}}{(\varepsilon_{n}-\varepsilon_{m})^{2}}({\sf\Lambda}_{i})_{nm}({\sf\Lambda}_{j})_{mn}\partial_{x_{i}}V\partial_{x_{j}}V+\mathcal{O}(\partial_{x}^{3}).$
The details of the derivation are described in appendix K.
3.2.3 Solutions to the kinetic equation and current (expansion in small potential)
eq. 116 is then solved by letting $f=f^{r}+\delta f$ and solving for $\delta f$ perturbatively in powers of $e$, $\delta f=\delta f_{e=1}+\delta f_{e=2}+\cdots$ where $\delta f_{e=i}\sim e^{i}$ for $i=1,2,\ldots$. If we define the Fourier transform, $\delta f_{e=i}(x,k,t)=\int_{q,\omega}\widehat{\delta f}_{e=i}(q,k,\omega)e^{iq\cdot x-i\omega t}$, we find
$\widehat{\delta f}_{e=1}(q,k,\omega)$ $=\frac{(n_{F}(k+q/2)-n_{F}(k-q/2))\,(1-{\sf g}_{jl}q_{j}q_{l})}{\hbar\omega-\varepsilon(k+q/2)+\varepsilon(k-q/2)+i\hbar\tau^{-1}}e\widehat{V}(q,\omega)+\mathcal{O}(q^{3}),$ (118)
$\widehat{\delta f}_{e=2}(q,k,\omega)$ $=e\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\frac{i{\sf\Omega}_{ij}q_{i}q_{j}^{\prime}+\hbar(\omega^{\prime}-\vec{v}\cdot{\vec{q}\,}^{\prime}){\sf t}_{ij}q_{i}q_{j}^{\prime}+\frac{1}{2}\partial_{k_{i}}{\sf g}_{jl}(q^{\prime\prime}_{i}q^{\prime}_{j}q^{\prime}_{l}+q_{i}^{\prime}q_{j}^{\prime\prime}q_{l}^{\prime\prime})}{\hbar(\omega-\vec{v}\cdot\vec{q}+i\tau^{-1})}\widehat{V}(q^{\prime},\omega^{\prime})\widehat{\delta f}_{e=1}(q^{\prime\prime},k,\omega^{\prime\prime})$ (119)
$+e\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\frac{1-{\sf g}_{ij}q_{j}q^{\prime}_{l}}{\hbar(\omega-\vec{v}\cdot\vec{q}+i\tau^{-1})}\widehat{V}(q^{\prime},\omega^{\prime})\left(\widehat{\delta f}_{e=1}\left(q^{\prime\prime},k+\frac{q^{\prime}}{2},\omega^{\prime\prime}\right)-\widehat{\delta f}_{e=1}\left(q^{\prime\prime},k-\frac{q^{\prime}}{2},\omega^{\prime\prime}\right)\right)$
$-e^{2}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\frac{n_{F}^{\prime}v_{i}{\sf t}_{jl}q_{i}q^{\prime\prime}_{j}q^{\prime\prime}_{l}}{\hbar(\omega-\vec{v}\cdot\vec{q}+i\tau^{-1})}\widehat{V}(q^{\prime},\omega^{\prime})\widehat{V}(q^{\prime\prime},\omega^{\prime\prime})-\frac{\omega-\vec{v}\cdot\vec{q}}{\omega-\vec{v}\cdot\vec{q}+i\tau^{-1}}\widehat{f}^{r}_{e=2}(q,k,\omega)+\mathcal{O}(q^{3})$
where the primed integral indicates an implicit delta function $\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}=\int_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}(2\pi)^{d+1}\delta^{d}(q-q^{\prime}-q^{\prime\prime})\delta(\omega-\omega^{\prime}-\omega^{\prime\prime})$, and $\widehat{V}$, $\widehat{f}^{r}_{e=2}$ are the Fourier transforms of $V$ and the $e^{2}$ term of $f^{r}$ respectively. For $\omega+i\hbar\tau^{-1}\neq 0$, the expressions above are valid to order $q^{4}$, but in the limit $\omega+i\hbar\tau^{-1}\rightarrow 0$, we find they are valid to order $q^{3}$. The solutions above can be used to calculate the electric current in (47). The $e$ contributions to the current vanish. The $e^{2}$ and $e^{3}$ contributions to the current are
$\widehat{J}_{i,e=2}(q,\omega)=$ $-e\int_{k}\mathrm{tr}\left[\widehat{\delta f}_{e=1}(q,k,\omega)\left(v_{i}-ie^{-1}{\sf m}_{ij}^{\textrm{orb}}q_{j}+\frac{1}{12\hbar}\left(\frac{1}{2}\partial_{k_{i}k_{j}k_{l}}^{3}\varepsilon-2(\partial_{k_{i}}{\sf c}_{jl}-\partial_{k_{j}}{\sf c}_{il}+2({\sf c}_{jl;i}-{\sf c}_{il;j}))\right)q_{j}q_{l}\right)\right]$
$-\frac{e^{2}}{\hbar}\int_{k}\mathrm{tr}[n_{F}(i{\sf\Omega}_{ij}q_{j}+\hbar(v_{i}{\sf t}_{jl}-v_{j}{\sf t}_{il})q_{j}q_{l}+(\partial_{k_{l}}{\sf g}_{ij}-{\sf s}_{ij;l})q_{j}q_{l})]\widehat{V}(q,\omega)+\mathcal{O}(q^{3}),$ (120)
$\widehat{J}_{i,e=3}(q,\omega)=$ $-e\int_{k}\mathrm{tr}\left[\widehat{\delta f}_{e=2}(q,k,\omega)\left(v_{i}-ie^{-1}{\sf m}_{ij}^{\textrm{orb}}q_{j}+\frac{1}{12\hbar}\left(\frac{1}{2}\partial_{k_{i}k_{j}k_{l}}^{3}\varepsilon-2(\partial_{k_{i}}{\sf c}_{jl}-\partial_{k_{j}}{\sf c}_{il}+2({\sf c}_{jl;i}-{\sf c}_{il;j}))\right)q_{j}q_{l}\right)\right]$
$-\frac{e^{2}}{\hbar}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\int_{k}\widehat{V}(q^{\prime},\omega^{\prime})\mathrm{tr}\left[\widehat{\delta f}_{e=1}(q^{\prime\prime},k,\omega^{\prime\prime})\left(i{\sf\Omega}_{ij}q^{\prime}_{j}-\hbar v_{j}{\sf t}_{il}q_{j}^{\prime}q_{l}^{\prime}-\frac{1}{2}\partial_{k_{i}}{\sf g}_{jl}q^{\prime\prime}_{j}q^{\prime}_{l}+\frac{1}{2}(\partial_{k_{l}}{\sf g}_{ij}+\partial_{k_{j}}{\sf g}_{il}-2{\sf s}_{ij;l})q_{j}q_{l}^{\prime}\right)\right]$
$-\frac{e^{3}}{\hbar}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\int_{k}\widehat{V}(q^{\prime},\omega^{\prime})\widehat{V}(q^{\prime\prime},\omega^{\prime\prime})\mathrm{tr}\left[n_{F}\left(\frac{1}{2}\partial_{k_{i}}{\sf t}_{jl}-\partial_{k_{j}}{\sf t}_{il}\right)\right]q_{j}^{\prime}q_{l}^{\prime\prime}-e\int_{k}\mathrm{tr}[\widehat{f}^{r}_{e=2}(q,k,\omega)v_{i}]+\mathcal{O}(q^{3}).$ (121)
Here, $\widehat{J}_{i,e=2},\widehat{J}_{i,e=3}$ describe linear and second-order electrical responses respectively, to second-order in $q$.
3.2.4 $v_{i}q_{i}\ll|\omega+i\tau^{-1}|$ limit
Let us now consider the limit in which $v_{i}q_{i},v_{i}q_{i}^{\prime},v_{i}q_{i}^{\prime\prime}\ll\lvert{\omega+i\tau^{-1}}\rvert$, so that we can expand the denominators of $\widehat{\delta f}_{e=i}$ in powers of $q$. In this limit $\widehat{J}_{i,e=2}(q,\omega)$ becomes
$\widehat{J}_{i,e=2}(q,\omega)=$ $\frac{e^{2}}{\hbar}\left(\frac{\hbar\tau\langle v_{i}v_{j}\rangle_{\textrm{FS}}}{1-i\omega\tau}-\sum_{n=1}^{N}\int_{k}n_{F,n}{\sf\Omega}_{n,ij}\right)iq_{j}\widehat{V}(q,\omega)+e^{2}\left(\frac{\tau^{2}\langle v_{i}v_{j}v_{l}\rangle_{\textrm{FS}}}{(1-i\omega\tau)^{2}}-\langle v_{l}{\sf g}_{ij}\rangle_{\textrm{FS}}\,\right)q_{j}q_{l}\widehat{V}(q,\omega)$ (122)
$+\frac{e^{2}}{\hbar}\left(\frac{\hbar\tau/e}{1-i\omega\tau}\langle v_{l}{\sf m}^{\textrm{orb}}_{ij}\rangle_{\textrm{FS}}+\sum_{n=1}^{N}\int_{k}n_{F,n}(\hbar(v_{n,j}{\sf t}_{n,il}-v_{n,i}{\sf t}_{n,jl})+{\sf s}_{n,ij;l})\right)q_{j}q_{l}\widehat{V}(q,\omega)+\mathcal{O}(q^{3}),$
where we defined $\langle d\rangle_{\textrm{FS}}\equiv\sum_{n=1}^{N}\int_{k}n_{F,n}^{\prime}d_{n}$ as an integral over the Fermi surface. The first two terms are the familiar Drude term ($\frac{e^{2}}{\hbar}\frac{\hbar\tau\langle v_{i}v_{j}\rangle_{\textrm{FS}}}{1-i\omega\tau}$) and Hall effect term ($-\frac{e^{2}}{\hbar}\sum_{n=1}^{N}\int_{k}n_{F,n}{\sf\Omega}_{n,ij}$) respectively. The other terms describe the electric current linear response to a non-uniform electric field. Notice that all the terms in the last line are divergenceless terms that describe edge currents and they are also multi-band quantities (meaning their calculation requires knowledge of the wavefunctions of multiple bands) in contrast to the terms in the first line which are all single-band quantities (whose calculation only requires knowledge of the wavefunction of a single band). Also note that a few terms in (122) remain nonzero even when the Fermi energy is in a band gap at $T=0$. This does not however contradict what we expect for a band insulator since those terms which are nonzero are divergenceless currents and conserve the local charge distribution.
Among the “divergenceful” terms is the contribution from the quantum metric ($e\langle v_{l}{\sf g}_{ij}\rangle_{\rm FS}$). Physically, this contribution comes from the electric quadrupole moment of the Bloch states which experiences a net force in a non-uniform electric field [16]. Rewriting this term as an integral over the Fermi sea, it can also be expressed as a contribution from the quantum-metric dipole (QMD), see (7).
A similar term was previously discussed in Ref. [16] by Lapa and Hughes. There, the authors generalized the semiclassical wave packet formalism of Ref. [6] by introducing a correction to the wave packet energy that comes from the electric quadrupole moment coupling to a non-uniform component of an electric field. They then derived the corrected equations of motion of the wave packet and solved the Boltzmann equation to find the contribution to the current $\widehat{J}_{\textrm{geom},i}\sim\frac{e^{2}}{2}\langle v_{i}{\sf g}_{jl}\rangle_{\textrm{FS}}q_{j}q_{l}\widehat{V}$, (rewritten to match our notation). Contrasting this with our result, we see the numerical prefactor as well as the permutation of the indices are different. [^3] Note that the choice of $f^{r}$ does not affect the quantum metric term which comes from an intrinsic (i.e. not collision-integral related) mechanism.. Experimental detection of such a non-uniform response has been proposed using nonreciprocal directional dichroism in Ref. [17], although the prediction there also differs from ours due to the above discrepancy.
Similarly, $\widehat{J}_{i,e=3}(q,\omega)$ becomes
$\widehat{J}_{i,e=3}(q,\omega)=$ $-\frac{e^{3}}{\hbar}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\bigg{(}\frac{\tau^{2}\langle v_{l}\partial_{k_{j}}v_{i}\rangle_{\textrm{FS}}}{(1-i\omega\tau)(1-i\omega^{\prime\prime}\tau)}+\frac{\tau\langle v_{l}{\sf\Omega}_{ij}\rangle_{\textrm{FS}}}{1-i\omega^{\prime\prime}\tau}-\frac{\hbar}{2}\langle v_{i}{\sf t}_{jl}\rangle_{\textrm{FS}}+\hbar\langle v_{j}{\sf t}_{il}\rangle_{\textrm{FS}}$ (123)
$\hskip 30.0pt-\frac{1}{1-i\omega\tau}\sum_{\begin{subarray}{c}n,m=1\\ n\neq m\end{subarray}}^{N}\int_{k}n_{F,n}\partial_{k_{i}}(\varepsilon_{n}-\varepsilon_{m})^{-1}({\sf\Lambda}_{j})_{nm}({\sf\Lambda}_{l})_{mn}\bigg{)}q^{\prime}_{j}q^{\prime\prime}_{l}\widehat{V}(q^{\prime},\omega^{\prime})\widehat{V}(q^{\prime\prime},\omega^{\prime\prime})+\mathcal{O}(q^{3}).$
Note that the second term of eq. 123, i.e. $-\frac{e^{2}}{\hbar}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\frac{\tau\langle v_{l}{\sf\Omega}_{ij}\rangle_{\textrm{FS}}}{1-i\omega^{\prime\prime}\tau}$, which is a transverse contribution to the current, is exactly the nonlinear Hall effect first proposed by Sodemann and Fu [21] that originates from the Berry curvature dipole.
If the Fermi energy is in a band gap at $T=0$, all terms except the last vanish, but unlike in the case of linear electrical response this finite term is not divergenceless. This occurs because the RTA violates local charge conservation in the system. The RTA couples the system to an external bath using a constant scattering rate, so even if the Fermi energy is in the band gap, a “divergenceful” current can still be generated. A pratical remedy for this issue is to take the collisionless limit $\tau^{-1}\rightarrow 0$ when the Fermi energy is in the band gap.
A general expression for second-order conductivity that ignored interband transitions was also studied in Ref. [18] by Kaplan et al. in which they used the RTA. Because their choice of $f^{r}$ in the RTA differs from ours and is $f^{r}=n_{F}(\varepsilon)$, which differs from ours, only the first line of eq. 123 here should be compared with the results of Ref. [18]. [Indeed the last line of eq. 123 comes from $f^{r}_{e=2}$ of eq. 117, which does not exist in Ref. [18].] Moreover, Kaplan et al. express their results as an integral over the Fermi sea, so we must we write our $\langle\cdots\rangle_{\textrm{FS}}$ terms by integrating by parts to compare them with their results. What we find is that only the first and last terms of the first line in eq. 123 match the results of Ref. [18] exactly. Contributions proportional to the second and third term are also present in the results of Ref. [18], but we find the numerical prefactors are different.
3.3 Fermi liquids with nontrivial band geometry
The quantum kinetic equation is powerful enough to calculate dynamical responses for quantities other than the current. Here we examine the dynamical density response functions, which in a metal display typical non-analytic frequency and momentum dependence at low energy. We recover the standard forms known for free electrons, and observe corrections arising from quantum geometry.
The real space fermion density $\rho(x,t)\equiv\langle\psi^{\dagger}(x,t)\psi(x,t)\rangle=\int_{k}\mathrm{tr}[{\sf F}(x,k,t)]$ can be calculated from the solution to the kinetic equation. Since the external potential $V(x,t)$ is coupled to $\psi^{\dagger}(x,t)\psi(x,t)$, the coefficients of the expansion of $\rho(x,t)$ in powers of $V$ are be density correlation functions.
Here, we consider the collisionless limit $\mathcal{I}[f]=0$, and assume $\omega\ll\omega_{\textrm{gap}}$ such that $\tilde{\mathcal{F}}$ is negligible. Then, $\rho(x,t)=\int_{k}\mathrm{tr}[f(x,k,t)]$, and we solve for $f$ using the kinetic equation eq. 116. Expanding $f$ perturbatively in powers of $e$ such that $f=n_{F}(\varepsilon)+f_{e=1}+f_{e=2}+\cdots$ where $f_{e=i}\sim e^{i}$ as in section 3.2.3, we find
$\widehat{f}_{e=1}(q,k,\omega)=$ $\frac{\hbar n_{F}^{\prime}\vec{v}\cdot\vec{q}\,(1-{\sf g}_{ij}q_{i}q_{j})}{\hbar\omega-\varepsilon(k+q/2)+\varepsilon(k-q/2)}e\widehat{V}$
$+n_{F}{\sf t}_{ij}q_{i}q_{j}e\widehat{V}+\mathcal{O}(q^{3}),$ (124)
and
$\widehat{f}_{e=2}(q,k,\omega)=$ $-\frac{e^{2}}{2}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}\mathscr{F}(\omega q;\omega^{\prime}q^{\prime};\omega^{\prime\prime}q^{\prime\prime};k)(1+{\sf g}_{ij}(q_{i}q^{\prime}_{j}+q^{\prime}_{i}q^{\prime\prime}_{j}+q^{\prime\prime}_{i}q_{j}))\widehat{V}(-q^{\prime},-\omega^{\prime})\widehat{V}(-q^{\prime\prime},-\omega^{\prime\prime})$ (125)
$-\frac{e^{2}}{2\hbar}\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}n_{F}^{\prime}\Bigg{[}\left(\frac{\vec{v}\cdot{\vec{q}\,}^{\prime}}{(\omega^{\prime}-\vec{v}\cdot{\vec{q}\,}^{\prime})(\omega-\vec{v}\cdot\vec{q})}-\frac{\vec{v}\cdot{\vec{q}\,}^{\prime\prime}}{(\omega^{\prime\prime}-\vec{v}\cdot{\vec{q}\,}^{\prime\prime})(\omega-\vec{v}\cdot\vec{q})}\right)(\partial_{k_{l}}{\sf g}_{ij}q^{\prime\prime}_{i}q^{\prime\prime}_{j}q^{\prime}_{l}+i{\sf\Omega}_{ij}q^{\prime\prime}_{i}q^{\prime}_{j})$
$\hskip 50.0pt+\left(\frac{q_{i}q^{\prime}_{j}q^{\prime\prime}_{l}}{\omega^{\prime\prime}-\vec{v}\cdot{\vec{q}\,}^{\prime\prime}}-\frac{q^{\prime}_{i}q^{\prime\prime}_{j}q^{\prime}_{l}}{\omega-\vec{v}\cdot\vec{q}}\right)2\hbar v_{l}{\sf t}_{ij}\Bigg{]}\widehat{V}(-q^{\prime},-\omega^{\prime})\widehat{V}(-q^{\prime\prime},-\omega^{\prime\prime})+\mathcal{O}(q^{3}),$
where $\widehat{f}_{e=i}$ is the Fourier transform of $f_{e=i}$ and $n_{F}=n_{F}(\varepsilon)$. The primed integrals above indicate an implicit delta function $\int^{\prime}_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}=\int_{\begin{subarray}{c}q^{\prime}q^{\prime\prime}\\ \omega^{\prime}\omega^{\prime\prime}\end{subarray}}(2\pi)^{d+1}\delta^{d}(q+q^{\prime}+q^{\prime\prime})\delta(\omega+\omega^{\prime}+\omega^{\prime\prime})$. Note that the convention for the signs in the delta function differ from the convention used in the previous subsection. In addition, $\mathscr{F}(\omega,q;\omega^{\prime}q^{\prime};\omega^{\prime\prime}q^{\prime\prime})$ is the integrand of the three-point density correlation function in the case of trivial band geometry which is defined as
$\mathscr{F}(\omega,q;\omega^{\prime},q^{\prime};\omega^{\prime\prime},q^{\prime\prime};k)=-\sum_{r=L,R}\sum_{i=1}^{3}n_{F}(k+Q_{r,i})\prod_{j\neq i}\frac{1}{\hbar(\Omega_{r,j}-\Omega_{r,i})-\varepsilon(k+Q_{r,j})+\varepsilon(k+Q_{r,i})},$ (126)
where $r=L,R$ labels the two contributing fermion loop diagrams and $(\Omega_{R,1},\Omega_{R,2},\Omega_{R,3})=(\omega,-\omega^{\prime\prime},0)$, $(Q_{R,1},Q_{R,2},Q_{R,3})=(q/2,(q^{\prime}-q^{\prime\prime})/2,-q/2)$ and $(\Omega_{L,1},\Omega_{L,2},\Omega_{L,3})=(\omega,-\omega^{\prime},0)$, $(Q_{L,1},Q_{L,2},Q_{L,3})=(q/2,-(q^{\prime}-q^{\prime\prime})/2,-q/2)$.
We can then compute the $n$-point density correlation functions $C^{(n)}(x_{1},t_{1},\ldots,x_{n},t_{n})=\langle\left(\psi^{\dagger}\psi\right)(x_{1},t_{1})\cdots\left(\psi^{\dagger}\psi\right)(x_{n},t_{n})\rangle$ for $n=2,3$. By differentiating (124) and (125) with respect to $\widehat{V}$, we obtain
$\widehat{C}^{(2)}(\omega,q)$ $=-\mathrm{i}e^{-1}\hbar\int_{k}\frac{\delta\widehat{f}_{e=1}(q,k,\omega)}{\delta\widehat{V}(q,\omega)}\approx-\mathrm{i}\int_{k}\left[\frac{\hbar^{2}n_{F}^{\prime}\vec{v}\cdot\vec{q}\,(1-{\sf g}_{ij}(k)q_{i}q_{j})}{\hbar\omega-\varepsilon(k+q/2)+\varepsilon(k-q/2)}+\hbar n_{F}{\sf t}_{ij}(k)q_{i}q_{j}\right],$ (127a)
$\widehat{C}^{(3)}(\omega^{\prime},q^{\prime},\omega^{\prime\prime},q^{\prime\prime})$ $=-e^{-2}\hbar^{2}\int_{k}\frac{\delta^{2}\widehat{f}_{e=2}(q,k,\omega)}{\delta\widehat{V}(q^{\prime},\omega^{\prime})\delta\widehat{V}(q^{\prime\prime},\omega^{\prime\prime})}$
$\approx\hbar^{2}\int_{k}\mathscr{F}(-\omega^{\prime}-\omega^{\prime\prime},-q^{\prime}-q^{\prime\prime};\omega^{\prime}q^{\prime};\omega^{\prime\prime}q^{\prime\prime};k)(1-{\sf g}_{ij}(k)(q^{\prime}_{i}q^{\prime}_{j}+q^{\prime}_{i}q^{\prime\prime}_{j}+q^{\prime\prime}_{i}q^{\prime\prime}_{j}))$ (127b)
$\quad-\hbar\int_{k}n_{F}^{\prime}\Bigg{[}\left(\frac{\vec{v}\cdot{\vec{q}\,}^{\prime}}{(\omega^{\prime}-\vec{v}\cdot{\vec{q}\,}^{\prime})(\omega^{\prime}+\omega^{\prime\prime}-\vec{v}\cdot(\vec{q}^{\prime}+\vec{q}^{\prime\prime}))}-\frac{\vec{v}\cdot{\vec{q}\,}^{\prime\prime}}{(\omega^{\prime\prime}-\vec{v}\cdot{\vec{q}\,}^{\prime\prime})(\omega^{\prime}+\omega^{\prime\prime}-\vec{v}\cdot(\vec{q}^{\prime}+\vec{q}^{\prime\prime}))}\right)$
×$\quad\quad\times\left(\partial_{k_{l}}{\sf g}_{ij}(k)q^{\prime\prime}_{i}q^{\prime\prime}_{j}q^{\prime}_{l}+i{\sf\Omega}_{ij}(k)q^{\prime\prime}_{i}q^{\prime}_{j}\right)+\left(\frac{(q^{\prime}_{i}+q^{\prime\prime}_{i})q^{\prime}_{j}q^{\prime\prime}_{l}}{\omega^{\prime\prime}-\vec{v}\cdot{\vec{q}\,}^{\prime\prime}}-\frac{q^{\prime}_{i}q^{\prime\prime}_{j}q^{\prime}_{l}}{\omega^{\prime}+\omega^{\prime\prime}-\vec{v}\cdot(\vec{q}^{\prime}+\vec{q}^{\prime\prime})}\right)2\hbar v_{l}{\sf t}_{ij}(k)\Bigg{]},$
where $\widehat{C}^{(n)}$ is the Fourier transform of $C^{(n)}$. The static structure factor can be obtained from the two-point density correlation function through $S_{q}\equiv\mathrm{Re}~{}\int\frac{\mathrm{d}\omega}{2\pi}\widehat{C}^{(2)}(\omega,q)$. Plugging in (127a) into this expression, we find, for $d=2$,
$\begin{split}S_{q}^{\mathrm{intra}}=&\mathrm{Im}~{}\int\frac{\mathrm{d}\omega}{2\pi}\int_{k}\frac{\hbar n_{F}^{\prime}\vec{v}\cdot\vec{q}\,(1-{\sf g}_{ij}(k)q_{i}q_{j})}{\omega-\vec{v}\cdot\vec{q}+\mathrm{i}\,\mathrm{sgn}(\omega)0^{+}}\\ =&\frac{\hbar^{2}k_{\mathrm{F}}}{2\pi^{2}}\left(|\vec{q}\,|-\frac{1}{4}\int_{0}^{2\pi}\mathrm{d}\theta~{}|\vec{n}\cdot\vec{q}\,|{\sf g}_{ij,\mathrm{F}}(\theta)q_{i}q_{j}\right),\end{split}$ (128)
where $k_{\mathrm{F}}$ is the Fermi wavevector and ${\sf g}_{ij,\mathrm{F}}(\theta)={\sf g}_{ij}(k=k_{\mathrm{F}},\theta)$. Notice that, as we assumed $\omega\ll\omega_{\mathrm{gap}}$, (128) only involves intraband contributions. In addition to the leading term, $\propto q$ ($\frac{\hbar^{2}k_{\rm F}}{2\pi^{2}}|\vec{q}|$), which is the single-band contribution to the Fermi gas, the quantum metric contributes at cubic order, $q^{3}$, to $S_{q}^{\mathrm{intra}}$ provided the $\theta$-integral in (128) is non-vanishing. These results can be compared to a Green’s function calculation of non-interacting multi-band fermions, which we carry out in appendix L.
4 Conclusion
4.1 Framework and comparison to other approaches
The theory of electrons in solids rests on the concept of energy bands and Bloch states. An essential element of the modern description of these bands is quantum geometry: not only the band energies but their wavefunctions enter into the dynamical response of electrons to various probes and forces, through quantities such as the Berry curvature and quantum metric. The latter quantities can be understood as related to the motion of electrons through a constrained Hilbert space of a single band or group of bands, rather than through the full Hilbert space. It is an axiom of the textbook semiclassical model [33] that the electronic response to weak field and forces, with slow spatial variations, can be described as independent motions within bands.
(i) The most rigorous approach to derive the connection of quantum geometric quantities to physical responses is through linear and non-linear response theory, i.e. generalized Kubo formulas. These formulas, based on time-dependent perturbation theory, often lead to simple expressions in terms of quantum geometric quantities, albeit through an obscure set of intermediate steps that make their origin opaque. (ii) The simpler expressions can, in some cases, be reproduced through a semiclassical wavepacket approach, which treats the single particle Schrödinger equation in a quasi-classical approximation, and then importing the latter data into a Boltzmann equation phenomenologically. This wavepacket approach [6] is intuitive but involves two successive approximations (the semi-classical one for the wavepacket, and the reduction of the full density matrix to a diagonal Boltzmann equation) which may be non-trivial to improve consistently together at higher orders in semi-classics. There, the quantum geometric contributions arise through projection to the band manifold. (iii) A third approach is the quantum kinetic equation, essentially exploiting the equation of motion for the full quantum density matrix. In a multi-band system the variables of this approach are matrix-valued, and include interband effects at the same level as intraband ones. This approach is fully rigorous, but the simplicity of the semiclassical limit, i.e. the existence of a description in terms of intraband motion, is, like in the Kubo approach, hidden.
In this paper, we have begun with the full matrix quantum kinetic equation and systematically derived a band-diagonal one to second order in the semi-classical expansion, which can be regarded as a formal expansion in $\hbar$, but is practically controlled as an expansion in spatial gradients. This is one order beyond the standard approach—first order leads to Berry curvature effects but not those of the quantum metric. [^4] One may wonder why they appear at different orders even if they can be united in a QGT. This is because at the leading order, the Berry curvature is linear in the Berry connection but the quantum metric is not, which itself comes from cancellations due to the fact that the Moyal product is antisymmetric but the quantum metric is symmetric. The diagonal dynamics enjoys a reduction of degrees of freedom: for an $N$-band system, it is expressed in terms of $N$ phase space densities rather than $N^{2}$ matrix elements. While this offers some computational gain, the main advantage is conceptual: one sees directly how all physics can be systematically expressed in terms of band-diagonal quantities, with quantum geometry emerging naturally through the diagonalization.
A recent preprint by Mitscherling et al. takes a similar approach in which they define a matrix-valued Wigner distribution function to study transport [34]. While their work is similar in spirit to ours, their formalism is developed for lattices and they do not include the effects of band geometry.
4.2 Is semi-classical dynamics purely band geometric?
A subtlety of band-projection is that a gauge freedom is introduced, corresponding to the freedom to redefine the phase of Bloch functions. In the semi-classical approach, the associated phase can be both position and momentum dependent, so that the gauge freedom resides in the full semi-classical phase space. Physical quantities must be gauge invariant. They may or may not be band-geometric in the single band sense. The Berry curvature and quantum metric are band geometric because they can be obtained from the wavefunctions of a single band only and do not require information about other bands or the Hamiltonian. The orbital magnetization is a quantity which is gauge invariant but not band geometric: to obtain it one needs to know the Hamiltonian or the full set of wavefunctions of all the bands. Many results in the literature involve quantities like the orbital magnetization which are not single-band geometric. Another example is the “Berry connection polarizability” also known as the “band normalized quantum metric” derived in Refs. [15, 19, 20, 18] as a contribution to non-linear conductivity. This suggests that the axiom of independent band dynamics is violated by some semiclassical corrections.
A remarkable feature of our results is that, when the Hamiltonian lacks explicit time dependence, the diagonal quantum kinetic equation can be expressed in a purely band-geometric form. In this limit, (33) becomes
$\partial_{t}f=$ $\frac{1}{i\hbar}[h\,\overset{\star}{,}\,f]-\partial_{\alpha}\left[\omega^{\alpha\beta}f\left(\hbar\omega^{\sigma\lambda}\partial_{\sigma}h\Omega_{\lambda\beta}\left(1-\frac{\hbar}{2}\omega^{\mu\nu}\Omega_{\mu\nu}\right)+\frac{\hbar^{2}}{2}\omega^{\sigma\lambda}\omega^{\mu\nu}\partial^{2}_{\sigma\mu}h\partial_{\beta}g_{\nu\lambda}\right)\right]$ (129)
$-\partial^{2}_{\alpha\sigma}\left[\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}f\left(\frac{1}{2}\omega^{\mu\nu}\partial_{\mu}h\partial_{\nu}g_{\beta\lambda}+\omega^{\mu\nu}\partial^{2}_{\mu\beta}hg_{\nu\lambda}\right)\right]+\mathcal{O}(\hbar^{3}).$
One observes that this equation involves only explicitly single-band-geometric quantities: the band energy $h$, the Berry curvature $\Omega_{\mu\nu}$ and the quantum metric $g_{\mu\nu}$.
This may be surprising given that it is well-known that multi-band quantities such as orbital magnetization and the quantum metric dipole do appear in physical responses. The resolution is that the fully band-diagonal form of eq. 129 appears only when quantities are expressed in terms of semi-classically corrected bands, which are the ones defined through Moyal diagonalization. When the kinetic equation is expressed in terms of quantities defined through the uncorrected bands (at zeroth order in $\hbar$), as worked out in sections 2.7 and 3 then the multi-band quantities appear, fully consistent with known results. Thus, our formulation gives the physical interpretation that the electronic dynamics (at least to second order in semiclassics) can be considered fully band-diagonal and intrinsically band-geometric, provided that the suitably renormalized bands are defined, and that the “magnetization-like” quantities which are not single-band in nature are only symptoms of expressing the renormalized bands in terms of bare ones.
4.3 Ramifications of the gradient expansion
The approach of this paper is based on the semi-classical $\hbar$ expansion, which amounts to an expansion in the smallness of spatial and temporal gradients. Thus it is clearly limited to situations in which the Hamiltonian varies slowly in real space and in time. The latter limits us to low frequencies, and generally ac response when external frequencies are comparable to or exceed interband transitions is not accessible. [^5] The extension to “fast” time variations/arbitrary frequencies is feasible, and not a fundamental limitation of the method. It would require solving for the full off-diagonal components of the distribution function $\tilde{\mathcal{F}}$, see eqs. 32 and 33 and section 3.2.1. This means we cannot address in a general way quantities like the shift current. There are, however, advantages to the present approach. First, in the semiclassical regime, the reduction to diagonal form is completely general, and eq. 129 and its extensions can be applied to arbitrary multi-band problems. Second, the approach allows treatment of general inhomogeneities in space. This allows a calculation of momentum dependent response functions, or non-equilibrium solutions for arbitrary spatial geometries.
By explicitly tracking the spatial dependence of physical quantities, the present approach directly confronts some of the subtle issues in transport theory. One of the basic such issues relates to the definition of current. A typical procedure to define a current is to extract it from the continuity equation. This defines a “transport current”: the current so-defined is guaranteed to describe the time dependence of the total charge in a closed region. That is, one takes a conserved quantity $Q$ and finds a local operator $\rho(x)$ such that $Q=\int_{x}\rho(x)$. Then one calculates $\partial_{t}\rho=i[H,\rho]$ which is unambiguous for a given Hamiltonian. One then finds a current $J_{i}(x)$ such that $\partial_{t}\rho=-\partial_{x_{i}}J_{i}$. However, this procedure has two ambiguities: the choice for $\rho(x)$ is not unique, and neither is the choice of $J_{i}$. The first ambiguity amounts to the freedom to shift $\rho\rightarrow\rho+\partial_{x_{i}}D_{i}$, where $D_{i}$ is a “dipole density”. If one makes this change to the definition of the density, then the definition of the current must change by $J_{i}\rightarrow J_{i}-\partial_{t}D_{i}$. This can be regarded as a polarization current. In a time-independent steady state this change does not affect the average $\langle J_{i}\rangle$ (because $\langle\partial_{t}D_{i}\rangle=0$). The polarization current will in general affect the non-zero frequency current response.
The second ambiguity is more treacherous. The (transport) current itself can be shifted by any curl, $J_{i}\rightarrow J_{i}+\partial_{x_{j}}M_{ij}$, where $M_{ij}$ is an anti-symmetric “magnetization density”, even when the definition of the density $\rho$ is unchanged. Due to the spatial derivative, this is not expected to change the zero momentum current in a translationally invariant system (for which Fourier decomposition is sensible), but it does affect non-zero wavevector currents, and it does affect the local currents when the system is not translationally invariant. The latter situation is important, for example, in the presence of even a constant temperature gradient, which leads to subtleties in calculations of the thermal Hall effect. However, by construction, the magnetization currents do not contribute to net transport through any closed surface or a surface that terminates outside the sample. This means that while they can, and do. affect calculations of the conductivity, the magnetization currents do not affect the conductance. In the formalism of this paper, it is possible to calculate a true conductance by treating a finite sample and obtaining the integrated current across a cross-section. We emphasize that the “Biot-Savart current” density, which is defined as the one that appears in Maxwell’s equations, is unambiguous. This is the one we derived in section 2.8, and has been used in this work.
4.4 Applications
In this paper, we applied our formalism to the case of a translationally-invariant Hamiltonian supplemented by a non-uniform electric field. We derived a form of the equilibrium current and also calculated the non-equilibrium current for the case of a time-dependent electric field. Our results can be compared to several others in the literature. We were inspired by Ref. [16], in which Lapa and Hughes generalized the wavepacket formalism to incorporate the effects of a non-uniform electric field and identified a quantum metric contribution to finite-momentum conductivity. However, using our formalism, we also find a quantum metric contribution but one that is distinct from theirs. Our results also reproduce the non-linear Hall effect of Ref. [21], but we find some discrepancies with our results and the quantum geometric non-linear conductivity of Ref. [18]. In addition to these transport quantities, we also obtain the dynamical density-density response functions. The two-point correlator has the standard (non-interacting) Fermi-liquid form, corrected by a higher order in momentum factor from the quantum metric. Our result evokes known quantum metric corrections for insulators discussed in Refs. [35, 36], but applies here at low frequencies.
4.5 Extensions
There are a number of interesting directions to extend this work. One could apply the formalism developed here to study thermal and thermo-electric currents and responses. It is particularly well-suited for this because a spatially inhomogeneous temperature can be explicitly treated (as can a non-thermal non-equilibrium state) without recourse to the “trick” of inserting an artificial gravitational field à la Luttinger, whose applicability beyond linear response is questionable. It would be natural to extend the current formalism to superonducting states, i.e. Bogoliubov-de Gennes Hamiltonians. An important direction is to include the effects of electron-electron or electron-phonon interactions. This could be accomplished via the Keldysh method, in which the semi-classical expansion is well-established for single band systems and can be best carried out by extending the Wigner transformation from just the space/momentum domain to include as well the time/frequency variables. The latter is essential to take into account frequency-dependent self-energy effects. We believe the techniques developed here should be a foundation for such a study in multi-band Fermi liquids. Another interesting direction is to extend our formalism to consider finite frequency responses beyond the low-frequency limit we considered in this work. This would involve solving for the off-diagonal components of the distribution function using eq. 32. This would allow us to study photovoltaic effects, shift and injection currents, and sum rules of optical conductivity which have been shown to be related to the integrated quantum metric [35, 37, 38]. Lastly, our formalism was developed assuming non-degenerate energy states but it would be of interest to extend our formalism to be fully compatible with degeneracies which generically occur in crystalline systems.
We hope that in the future this method can be used profitably in novel comparisons to experiments. Specifically, our formalism in phase space has a unique advantage of affording us both real space and momentum space resolution. It would then be interesting to try to apply this method to understand experimental measurements that probe the momentum structure of materials locally such as nano-angle-resolved photo-emission spectroscopy and interference patterns in scanning tunneling microscopy.
Acknowledgements
X.H. thanks Jing-Yuan Chen for interesting discussions. L.S. and L.B. thank Léo Mangeolle for a previous collaboration which led to this work. We especially thank Ren-Bo Wang and Kaixiang Su for many valuable discussions and collaboration at the early stage of this project. X.H. thanks the Kavli Institute for Theoretical Physics for hospitality; KITP is supported in part by the Heising-Simons Foundation, the Simons Foundation, and the National Science Foundation under Grant PHY-2309135. T.P. was funded by the Air Force Office of Scientific Research (AFOSR) (award FA9550-22-1-0432). L.S. was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 853116, acronym TRANSPORT). L.B. was supported by the NSF CMMT program under Grants No. DMR-2419871, and the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (Grant No. 651440).
Statement of author contributions
T.P. and X.H. contributed equally to this work, and carried out the majority of calculations. L.B. and L.S. conceptualized the project. All authors contributed to and checked analytic results, and took part in writing of the manuscript.
Appendix A Notations
  • non-diagonal quantities: sans-serif fonts, e.g. $\sf F,H$. We reserve the sans-serif fonts of other letters to distinguish for quantities defined in wave number. We also defined ${\sf P}_{n}=P_{n}^{(0)}$.
  • rotated gauge-dependent quantity : $\tilde{F},\tilde{H}$
  • rotated (non-diagonal) gauge-independent quantity : $F,H$
  • gauge-dependent diagonal quantities: lowercase, e.g. $\tilde{f},\tilde{h}$
  • off-diagonal quantities: $\tilde{\mathcal{F}}$
  • gauge-invariant diagonal quantities: $f,h$
  • Fourier transform with respect to the center of mass coordinate, $x\rightarrow q$: wide hat, e.g. $\widehat{V}$
  • band indices: Roman letters $n,m,m^{\prime},\cdots$
  • phase space coordinate indices: Greek letters $\alpha,\beta,\lambda,\sigma,\cdots$
  • real space coordinate indices: Roman letters $i,j,\cdots$
  • Moyal algebra: $\omega^{\alpha\beta}$, and $\omega^{\alpha\beta}_{B}$ in the presence of a $B$ field.
  • Berry connection: $\Lambda_{\alpha}=-\mathrm{i}U^{\dagger}\star\partial_{\alpha}U$ and $A_{\alpha}=\operatorname{diag}(\Lambda_{\alpha})$
  • External vector potential: $\mathcal{A}_{i}$
  • Physical current: $J_{i}(x)$
  • $\hbar$ expansion: $h^{(n)}\sim\hbar^{n}$ such that $h=h^{(0)}+h^{(1)}+\cdots$
  • external potential strength expansion: $f_{e=n}$ such that $f=f_{e=0}+f_{e=1}+f_{e=2}+f_{e=3}+\cdots$
Note that the notations in this manuscript differ significantly from those in a manuscript by two of us, Ref. [5]. Moreover, beyond notations, it is noteworthy that (i) the convention for $\Lambda_{\alpha}$ is different in the two manuscripts (in Ref. [5], $\Lambda_{\alpha}=U^{\dagger}\star\partial_{\alpha}U$ while here $\Lambda_{\alpha}=-iU^{\dagger}\star\partial_{\alpha}U$) but the $A_{\alpha}$ agree ($A_{\alpha}={\rm Im}[{\rm diag}\Lambda_{\alpha}]$ in Ref. [5] and here $A_{\alpha}={\rm diag}\Lambda_{\alpha}$), (ii) the matrices in Ref. [5] are $2N\times 2N$ matrices (compared to the $N\times N$ matrices used here) because of the lack of particle conservation in Ref. [5].
Appendix B Moyal product identities
Here, we list two useful formulas that are repeatedly used in our formalism.
Identity 1       Assuming $A,C$ are matrix-valued functions and $b$ a scalar-valued function,
$A\star b\star C=$ $b\,(A\star C)-\frac{i\hbar}{2}\omega^{\alpha\beta}\partial_{\alpha}b\,(\partial_{\beta}A\star C-A\star\partial_{\beta}C)-\frac{\hbar^{2}}{8}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}b\,(\partial^{2}_{\beta\lambda}(A\star C)-4\partial_{\beta}A\star\partial_{\lambda}C)+\mathcal{O}(\hbar^{3}).$ (130)
Identity 2       Assuming $A,B$ are matrix-valued functions,
$\mathrm{tr}[A\star B]=\mathrm{tr}[B\star A]-i\hbar\omega^{\alpha\beta}\mathrm{tr}[\partial_{\alpha}B\star\partial_{\beta}A]-\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\mathrm{tr}[\partial^{2}_{\alpha\sigma}B\star\partial^{2}_{\beta\lambda}A]+\mathcal{O}(\hbar^{3}).$ (131)
Appendix C Defining $f,h$
Here, we define $f,h$ which are the gauge-invariant forms of $\tilde{f},\tilde{h}$ respectively. Firstly, we define $f$ such that it satisfies
${\rm N}=\int_{x,p}\mathrm{tr}[{\sf F}]=\int_{x,p}\mathrm{tr}[f].$ (132)
Note, $\int_{x,p}\mathrm{tr}[{\sf F}]=\int_{x,p}\mathrm{tr}[U\star\tilde{F}\star U^{\dagger}]=\int_{x,p}\mathrm{tr}[U\star\tilde{f}\star U^{\dagger}]+\int_{x,p}\mathrm{tr}[U\star\tilde{\mathcal{F}}\star U^{\dagger}]=\int_{x,p}\mathrm{tr}[U\star\tilde{f}\star U^{\dagger}]$. The last equality was obtained by using the fact that the terms in the trace can be cycled even in the presence of star products upon taking the phase-space integral. The last equality tells us that the off-diagonal components of the distribution function does not contribute to the total particle number. We can then rewrite eq. 132 as
$\int_{x,p}\mathrm{tr}[f]=\sum_{n}\int_{x,p}\mathrm{tr}[U\star p_{n}\tilde{f}_{n}\star U^{\dagger}]$ (133)
where $p_{n}$ is a diagonal matrix defined such that $(p_{n})_{ij}=\delta_{in}\delta_{jn}$. The simplest definition for $f$ that satisfies this condition is
$f_{n}\equiv\mathrm{tr}[U\star p_{n}\tilde{f}_{n}\star U^{\dagger}].$ (134)
Using eqs. 130 and 131, the right-hand side can be expanded to second-order in $\hbar$ to give us
$f_{n}=\tilde{f}_{n}+\hbar\omega^{\alpha\beta}\partial_{\alpha}(\tilde{f}_{n}A_{n\beta})+\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}(\tilde{f}_{n}(\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\})_{nn})+\mathcal{O}(\hbar^{3}),$ (135)
and inverting this relation gives us
$\tilde{f}_{n}=f_{n}-\hbar\omega^{\alpha\beta}\partial_{\alpha}(f_{n}(A_{n\beta}+\hbar\omega^{\sigma\lambda}A_{n\lambda}\partial_{\sigma}A_{n\beta}))+\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\left(f_{n}\left(A_{n\beta}A_{n\lambda}-\frac{1}{4}\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\}_{nn}\right)\right)+\mathcal{O}(\hbar^{3}).$ (136)
Next, let us consider the definition for $h$. We require that $h$ satisfies
${\rm E}=\frac{1}{2}\int_{x,p}\mathrm{tr}[{\sf H}\star{\sf F}+{\sf F}\star{\sf H}]=\int_{x,p}\mathrm{tr}[hf].$ (137)
The expectation value of the Hamiltonian can be rewritten in the following way.
$\langle\hat{H}\rangle=$ $\frac{1}{2}\int_{x,p}\mathrm{tr}[{\sf H}\star{\sf F}+{\sf F}\star{\sf H}]$
$=$ $\sum_{n}\frac{1}{2}\int_{x,p}\mathrm{tr}[U\star p_{n}(\tilde{h}_{n}\star\tilde{f}_{n}+\tilde{f}_{n}\star\tilde{h}_{n})\star U^{\dagger}]$
$=$ $\sum_{n}\int_{x,p}\left(\tilde{h}_{n}\tilde{f}_{n}+\hbar\omega^{\alpha\beta}\partial_{\alpha}(\tilde{h}_{n}\tilde{f}_{n}A_{n\beta})-\frac{\hbar^{2}}{8}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{h}_{n}\partial^{2}_{\beta\lambda}\tilde{f}_{n}+\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}(\tilde{h}_{n}\tilde{f}_{n}(\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\})_{nn})\right)+\mathcal{O}(\hbar^{3})$
$=$ $\sum_{n}\int_{x,p}f_{n}\left(\tilde{h}_{n}+\hbar\omega^{\alpha\beta}\partial_{\alpha}\tilde{h}_{n}(A_{n\beta}+\hbar\omega^{\sigma\lambda}A_{n\lambda}\partial_{\sigma}A_{n\beta})-\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{h}_{n}\left(\frac{1}{4}\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\}-A_{n\beta}A_{n\lambda}\right)\right)+\mathcal{O}(\hbar^{3})$
Like before, the off-diagonal components of the distribution function contribute nothing. The last line was obtained by expressing $\tilde{f}_{n}$ using $f_{n}$ and integrating by parts. The last line gives us the definition of $h_{n}$.
$h_{n}\equiv\tilde{h}_{n}+\hbar\omega^{\alpha\beta}\partial_{\alpha}\tilde{h}_{n}(A_{n\beta}+\hbar\omega^{\sigma\lambda}A_{n\lambda}\partial_{\sigma}A_{n\beta})-\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{h}_{n}\left(\frac{1}{4}\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\}_{nn}-A_{n\beta}A_{n\lambda}\right)+\mathcal{O}(\hbar^{3})$ (138)
Its inverse relation is
$\tilde{h}_{n}=h_{n}-\hbar\omega^{\alpha\beta}\partial_{\alpha}h_{n}A_{n\beta}+\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}h_{n}(\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\})_{nn})+\mathcal{O}(\hbar^{3}).$ (139)
Appendix D Diagonalization by flow equations
The diagonalization equation eq. 13 which we reproduce below
$\tilde{h}=U^{\dagger}\star{\sf H}\star U$ (140)
can be solved using what we call “flow equations.” The basic idea behind flow equations is to take derivatives with respect to $\hbar$. First, we differentiate the unitarity condition of $U$ eq. 12 to get
$T^{\dagger}+T+\frac{\mathrm{i}\hbar}{2}\omega^{\alpha\beta}\Lambda_{\alpha}\star\Lambda_{\beta}=0$ (141)
where we defined $T\equiv\hbar U^{\dagger}\star\partial_{\hbar}U$. The second term comes from the derivative acting on the $\hbar$ in the definition of the star product. eq. 141 completely determines the hermitian part of $T$ so we can express $T$ as
$T=Y-\frac{\mathrm{i}\hbar}{4}\omega^{\alpha\beta}\Lambda_{\alpha}\star\Lambda_{\beta}=Y+\frac{\hbar}{4}\omega^{\alpha\beta}\partial_{\alpha}\Lambda_{\beta}$ (142)
where $Y$ is the anti-Hermitian part of $T$, $Y^{\dagger}=-Y$. Next, by differentiating the diagonalization equation (140) with respect to $\hbar$, we obtain, what we refer to as, the flow equation of $\tilde{h}$
$\begin{split}\hbar\partial_{\hbar}\tilde{h}=&\tilde{h}^{(1)}+2\tilde{h}^{(2)}+\cdots=[\tilde{h}\,\overset{\star}{,}\,Y]+\frac{\hbar}{2}\omega^{\alpha\beta}\{\Lambda_{\alpha}\,\overset{\star}{,}\,\partial_{\beta}\tilde{h}\}+\frac{i\hbar}{4}\omega^{\alpha\beta}\{\Lambda_{\alpha}\,\overset{\star}{,}\,[\Lambda_{\beta}\,\overset{\star}{,}\,\tilde{h}]\}.\end{split}$ (143)
The first-order terms of the right-hand side gives us the first-order correction to $\tilde{h}^{(0)}$.
$\begin{split}\tilde{h}^{(1)}&=[\tilde{h}^{(0)}\,,Y^{(1)}]+\frac{\hbar}{2}\omega^{\alpha\beta}\{\Lambda_{\alpha}^{(0)}\,,\partial_{\beta}\tilde{h}^{(0)}\}\quad+\frac{i\hbar}{4}\omega^{\alpha\beta}\{\Lambda^{(0)}_{\alpha}\,,[\Lambda^{(0)}_{\beta}\,,\tilde{h}^{(0)}]\}.\end{split}$ (144)
Since $\tilde{h}^{(1)}$ is supposed to be a diagonal matrix, this imposes a constraint on $Y^{(0)}$ which determines its off-diagonal components.
$Y^{(1)}_{nm}$ $\equiv-\frac{\hbar}{2(\tilde{h}^{(0)}_{n}-\tilde{h}^{(0)}_{m})}\omega^{\alpha\beta}\bigg{(}\{\Lambda_{\alpha}^{(0)}\,,\partial_{\beta}\tilde{h}^{(0)}\}+\frac{i}{2}\{\Lambda^{(0)}_{\alpha}\,,[\Lambda^{(0)}_{\beta}\,,\tilde{h}^{(0)}]\}\bigg{)}_{nm}\,,\quad n\neq m.$ (145)
Therefore $\tilde{h}^{(1)}$ is
$\begin{split}\tilde{h}^{(1)}=&-\hbar\omega^{\alpha\beta}\partial_{\alpha}\tilde{h}^{(0)}A_{\beta}^{(0)}+\frac{i\hbar}{4}\omega^{\alpha\beta}\operatorname{diag}[\{\Lambda^{(0)}_{\alpha}\,,[\Lambda^{(0)}_{\beta}\,,\tilde{h}^{(0)}]\}],\end{split}$ (146)
and from eq. 27, the gauge-invariant form of this term is
$h^{(1)}=\frac{i\hbar}{4}\omega^{\alpha\beta}\operatorname{diag}[\{\Lambda^{(0)}_{\alpha}\,,[\Lambda^{(0)}_{\beta}\,,\tilde{h}^{(0)}]\}].$ (147)
This term is a contribution from the phase-space orbital magnetization. Note that the diagonal components of $Y^{(1)}$ does not affect the diagonalization procedure, and we can let $\operatorname{diag}(Y^{(1)})=0$ without consequence.
The second-order correction $\tilde{h}^{(2)}$ can be similarly calculated from the flow equation eq. 143 by taking first-order terms of the right-hand side. However, it is a tedious calculation. We first define $\tilde{w}$ as the last term in the flow equation eq. 143:
$\tilde{w}=\frac{i\hbar}{4}\omega^{\alpha\beta}\{\Lambda_{\alpha}\,\overset{\star}{,}\,[\Lambda_{\beta}\,\overset{\star}{,}\,\tilde{h}]\}$ (148)
Consider the $n$-th diagonal of $\tilde{w}$. We find that it can be expressed as
$\tilde{w}_{n}=$ $-\frac{i\hbar}{2}\omega^{\alpha\beta}\mathrm{tr}[\partial_{\alpha}P_{n}\star{\sf H}\star\partial_{\beta}P_{n}]+\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial_{\alpha}(h_{n}\partial_{\sigma}g_{n,\beta\lambda})+\frac{\hbar^{2}}{8}h_{n}(\omega^{\alpha\beta}\Omega_{n,\alpha\beta})^{2}$ (149)
$+\hbar\omega^{\alpha\beta}\partial_{\alpha}\tilde{w}_{n}A_{n,\beta}+\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial_{\alpha}h_{n}A_{n,\sigma}\partial_{\beta}A_{n,\lambda}+\mathcal{O}(\hbar^{3}).$
Notice that the first-order component $\tilde{w}_{n}^{(1)}$ comes solely from the first term on the right-hand side, and this term is defined using only gauge-invariant objects. Therefore, $\tilde{w}_{n}^{(1)}$ is gauge-invariant and it is exactly $h_{n}^{(1)}$. The two other terms in the first line of the right-hand side are also gauge-invariant. Therefore, we define
$w_{n}\equiv-\frac{i\hbar}{2}\omega^{\alpha\beta}\mathrm{tr}[\partial_{\alpha}P_{n}\star{\sf H}\star\partial_{\beta}P_{n}]$ (150)
and write
$\tilde{w}_{n}=w_{n}+\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial_{\alpha}(h_{n}\partial_{\sigma}g_{n,\beta\lambda})+\frac{\hbar^{2}}{8}h_{n}(\omega^{\alpha\beta}\Omega_{n,\alpha\beta})^{2}\\ +\hbar\omega^{\alpha\beta}\partial_{\alpha}w_{n}A_{n,\beta}+\frac{\hbar}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial_{\alpha}h_{n}A_{n,\sigma}\partial_{\beta}A_{n,\lambda}+\mathcal{O}(\hbar^{3})$ (151)
If we now take the second-order terms of the right-hand side of the flow equation eq. 143, we see that
$\tilde{h}_{n}^{(2)}=-\frac{\hbar}{2}\omega^{\alpha\beta}(\partial_{\alpha}\tilde{h}^{(1)}A_{n\beta}^{(0)}+\partial_{\alpha}\tilde{h}^{(0)}A_{n\beta}^{(1)})+\frac{1}{2}\tilde{w}_{n}^{(2)}$ (152)
Then using eq. 27, we find
$\begin{split}h^{(2)}=&\frac{1}{2}w^{(2)}+\frac{\hbar}{2}\omega^{\alpha\beta}\partial_{\alpha}h^{(0)}\left(A_{\beta}^{(1)}+\hbar\omega^{\sigma\lambda}\left(\partial_{\sigma}A_{\beta}^{(0)}A_{\lambda}^{(0)}+\frac{1}{2}A_{\sigma}^{(0)}\partial_{\beta}A_{\lambda}^{(0)}+\frac{1}{2}\partial_{\sigma}g_{\beta\lambda}^{(0)}\right)\right)\\ &+\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}h^{(0)}\left(\partial^{2}_{\alpha\sigma}g_{\beta\lambda}^{(0)}+\frac{1}{4}\Omega_{\alpha\beta}^{(0)}\Omega_{\sigma\lambda}^{(0)}\right)-\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}h^{(0)}g_{\beta\lambda}^{(0)}\end{split}$ (153)
Using the definition for $m_{n,\alpha\beta}$ given in eq. 40 we see that $w_{n}=\frac{\hbar}{2}\omega^{\alpha\beta}m_{n,\alpha\beta}-\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}(h_{n}\partial^{2}_{\alpha\sigma}g_{n,\beta\lambda}+2\partial_{\alpha}h_{n}\partial_{\sigma}g_{n,\beta\lambda})+\mathcal{O}(\hbar^{3})$. If we now substitute this into the equation above, we get eq. 42.
In order to evaluate the second-order correction $h^{(2)}$ we also need to calculate the first-order correction to the Berry connection and the projection operator. We can do this by calculating their flow equations.
$\hbar\partial_{\hbar}\Lambda_{\alpha}=\left[\Lambda_{\alpha}\,\overset{\star}{,}\,Y\right]-i\partial_{\alpha}Y+\frac{\hbar}{4}\omega^{\sigma\lambda}\left\{\Lambda_{\sigma}\,\overset{\star}{,}\,\partial_{\lambda}\Lambda_{\alpha}\right\},$ (154)
and
$\begin{split}&\hbar\partial_{\hbar}P_{n}=U\star\bigg{(}\left[Y\,\overset{\star}{,}\,p_{n}\right]+\frac{\hbar}{4}\omega^{\alpha\beta}\left\{\partial_{\alpha}\Lambda_{\beta}\,\overset{\star}{,}\,p_{n}\right\}+\frac{i\hbar}{2}\omega^{\alpha\beta}\Lambda_{\alpha}\star p_{n}\Lambda_{\beta}\bigg{)}\star U^{\dagger}\end{split}$ (155)
which are obtained by taking the derivative of eq. 19 and eq. 22 with respect to $\hbar$. From these flow equations we get
$\begin{split}A_{\alpha}^{(1)}=&\operatorname{diag}\left([\Lambda_{\alpha}^{(0)}\,,Y^{(1)}]\right)+\frac{\hbar}{4}\omega^{\sigma\lambda}\operatorname{diag}\left(\{\Lambda_{\sigma}^{(0)}\,,\partial_{\lambda}\Lambda_{\alpha}^{(0)}\}\right),\end{split}$ (156)
and
$\begin{split}&P_{n}^{(1)}=U^{(0)}\bigg{(}[Y^{(1)}\,,p_{n}]+\frac{\hbar}{4}\omega^{\alpha\beta}\{\partial_{\alpha}\Lambda^{(0)}_{\beta}\,,p_{n}\}+\frac{i\hbar}{2}\omega^{\alpha\beta}\Lambda_{\alpha}^{(0)}p_{n}\Lambda^{(0)}_{\beta}\bigg{)}U^{(0)\dagger}.\end{split}$ (157)
Appendix E Derivation of the equilibrium distribution
Here, we derive the form for the equilibrium distribution function $\tilde{f}_{\textrm{eq}}$. We begin by defining the imaginary-time Green’s function ${\sf G}_{nm}(x_{1},\tau_{1};x_{2},\tau_{2})\equiv-\langle T_{\tau}\psi_{n}(x_{1},\tau)\psi^{\dagger}_{m}(x_{2},\tau_{2})\rangle_{\beta}$, where $\langle\cdots\rangle_{\beta}$ denotes an average with respect to a grand canonical ensemble at temperature $T=\frac{1}{\beta k_{B}}$. The equilibrium density matrix in the position representation can then be expressed as
${\sf F}_{\textrm{eq},nm}(x_{1},x_{2})\equiv\langle\psi^{\dagger}_{m}(x_{1})\psi_{n}(x_{2})\rangle_{\beta}={\sf G}_{nm}(x_{1},\tau;x_{2},\tau+0^{+})=\frac{1}{\beta}\sum_{i\omega_{n}}{\sf G}_{nm}(x_{1},x_{2};\omega_{n})$ (158)
where ${\sf G}_{nm}(x_{1},x_{2};\omega_{n})$ is the Matsubara-frequency representation of the imaginary-time Green’s function. It is defined implicitly by the equations
$\int_{y}\left[(i\omega_{n}+\mu)\delta^{d}(x-y)-{\sf H}(x,y)\right]{\sf G}(y,x^{\prime};\omega_{n})=$ $\delta(x-x^{\prime})I_{N}$ (159)
$\int_{y}{\sf G}(x,y;\omega_{n})\left[(i\omega_{n}+\mu)\delta^{d}(y-x^{\prime})-{\sf H}(y,x^{\prime})\right]=$ $\delta(x-x^{\prime})I_{N}$ (160)
where we have suppressed matrix indices for brevity. Let’s now transform this using the Wigner transformation.
$\left[i\omega_{n}+\mu-{\sf H}(x,p)\right]\star{\sf G}(x,p;\omega_{n})=$ $I_{N}$ (161)
${\sf G}(x,p;\omega_{n})\star\left[i\omega_{n}+\mu-{\sf H}(x,p)\right]=$ $I_{N}$ (162)
We now assume that all functions are defined on phase space and drop the $(x,p)$ dependence of functions. The next step is to apply the transformation $U$ which diagonalizes ${\sf H}$.
$\left[i\omega_{n}+\mu-\tilde{h}\right]\star\tilde{G}=$ $I_{N}$ (163)
$\tilde{G}\star\left[i\omega_{n}+\mu-\tilde{h}\right]=$ $I_{N}$ (164)
where $\tilde{G}\equiv U^{\dagger}\star{\sf G}\star U$.
We can now solve for $\tilde{G}$ in powers of $\hbar$. Let $\tilde{G}=\tilde{G}^{(0)}+\tilde{G}^{(1)}+\tilde{G}^{(2)}+\cdots$ where $\tilde{G}^{(n)}\sim\hbar^{n}$. To lowest order we find
$\tilde{G}^{(0)}=\frac{1}{i\omega_{n}-h^{(0)}+\mu}.$ (165)
A straightforward calculation also leads to the first- and second-order corrections:
$\tilde{G}^{(1)}=\frac{\tilde{h}^{(1)}}{\left(i\omega_{n}-h^{(0)}+\mu\right)^{2}}$ (166)
$\begin{split}\tilde{G}^{(2)}=&\frac{\tilde{h}^{(2)}}{\left(i\omega_{n}-h^{(0)}+\mu\right)^{2}}+\frac{(\tilde{h}^{(1)})^{2}}{\left(i\omega_{n}-h^{(0)}+\mu\right)^{3}}-\frac{1}{8}\omega^{\alpha\beta}\omega^{\sigma\lambda}\frac{\partial^{2}_{\alpha\sigma}h^{(0)}\partial^{2}_{\beta\lambda}h^{(0)}}{\left(i\omega_{n}-h^{(0)}+\mu\right)^{3}}-\frac{1}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\frac{\partial^{2}_{\alpha\sigma}h^{(0)}\partial_{\beta}h^{(0)}\partial_{\lambda}h^{(0)}}{\left(i\omega_{n}-h^{(0)}+\mu\right)^{4}}\end{split}$ (167)
We have now solved for $\tilde{G}$ up to second-order in $\hbar$.
Going back to eq. 158, if we apply $U$ to this we get $\tilde{f}_{\textrm{eq}}=\beta^{-1}\sum_{i\omega_{n}}\tilde{G}(i\omega_{n})$. Therefore, now all we need to do is sum over the Matsubara frequencies. We use the formula
$\frac{1}{\beta}\sum_{i\omega_{n}}\frac{e^{i\omega_{n}0^{+}}}{(i\omega_{n}-h^{(0)}+\mu)^{k}}=\frac{1}{(k-1)!}n_{F}^{(k-1)}(h^{(0)}-\mu)$ (168)
where $n_{F}^{(k-1)}$ is the $k-1$-th derivative of $n_{F}(x)=(e^{\beta(x-\mu)}+1)^{-1}$ which is the Fermi-Dirac distribution function. We then find
$\tilde{f}_{\textrm{eq}}=n_{F}(\tilde{h})-\frac{\hbar^{2}}{16}n_{F}^{\prime\prime}(\tilde{h})\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{h}\partial^{2}_{\beta\lambda}\tilde{h}-\frac{\hbar^{2}}{24}n_{F}^{\prime\prime\prime}(\tilde{h})\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{h}\partial_{\beta}\tilde{h}\partial_{\lambda}\tilde{h}+\mathcal{O}(\hbar^{3})$ (169)
Notably, this function can also be derived from the Moyal-generalization of the Fermi-Dirac distributions. Specifically, we define
$\tilde{f}_{\mathrm{eq}}=n_{\mathrm{F}}(-\mu)+n_{\mathrm{F}}^{\prime}(-\mu)\tilde{h}+\frac{1}{2}n_{\mathrm{F}}^{\prime\prime}(-\mu)\tilde{h}\star\tilde{h}+\frac{1}{3!}n_{\mathrm{F}}^{\prime\prime\prime}(-\mu)\tilde{h}\star\tilde{h}\star\tilde{h}+\ldots,$ (170)
which reduces to $n_{\mathrm{F}}(h-\mu)$ if all the star products are the normal products. By properly arranging the infinite series, we find it is precisely (169) to the second order in $\hbar$.
Appendix F Derivation of the electric current eq. 46
We assume our system is coupled to an external $U(1)$ gauge field $\mathcal{A}_{\mu}(x)$. Then, the associated charge current is given by
$J_{\mu}(x)=-\frac{\delta\langle\hat{H}_{\mathcal{A}}\rangle}{\delta\mathcal{A}_{\mu}(x)}$ (171)
where $\hat{H}_{\mathcal{A}}$ is the Hamiltonian.
$\langle\hat{H}_{\mathcal{A}}\rangle=\int_{x,p}\mathrm{tr}[{\sf H}_{\mathcal{A}}(x,p){\sf F}(x,p)].$ (172)
Below we show that ${\sf H}_{\mathcal{A}}(x,p)$ can be expressed to second order in $\hbar$ as
${\sf H}_{\mathcal{A}}(x,p)={\sf H}(x,\pi)-\frac{e\hbar^{2}}{24}\frac{\partial^{2}\mathcal{A}_{\mu}(x)}{\partial x_{\nu}\partial x_{\gamma}}\frac{\partial^{3}{\sf H}(x,\pi)}{\partial\pi_{\mu}\partial\pi_{\nu}\partial\pi_{\gamma}}+\mathcal{O}(\hbar^{3})$ (173)
where ${\sf H}(x,p)$ is the Hamiltonian when $\mathcal{A}=0$ and $\pi_{\mu}(x,p)=p_{\mu}+e\mathcal{A}_{\mu}(x)$. Then, the $U(1)$ current is given by
$J_{\mu}(x)=-e\int_{p}\mathrm{tr}\left[\frac{\partial{\sf H}(x,\pi)}{\partial\pi_{\mu}}{\sf F}(x,p)\right]+\frac{e\hbar^{2}}{24}\frac{\partial^{2}}{\partial x_{\nu}\partial x_{\gamma}}\int_{p}\mathrm{tr}\left[\frac{\partial^{3}{\sf H}(x,\pi)}{\partial\pi_{\mu}\partial\pi_{\nu}\partial\pi_{\gamma}}{\sf F}\right]+\mathcal{O}(\hbar^{3}).$ (174)
We first prove eq. 173 by considering a separable Hamiltonian meaning $\hat{H}=H_{0}(\hat{p})+H_{1}(\hat{x})$. Then, we consider the more general case.
We assume spinless fermions and introduce the external field via minimal coupling. Then, $\hat{H}_{\mathcal{A}}=H_{0}(\hat{p}+e\mathcal{A}(\hat{x}))+H_{1}(\hat{x})$. The second term $H_{1}(\hat{x})$ does not contribute to the current, so we ignore it. In general $H_{0}(p)$ can be expressed as
$H_{0}(\hat{p})=\sum_{n=1}^{\infty}\frac{1}{n!}G_{\alpha_{1}\cdots\alpha_{n}}\,\hat{p}_{\alpha_{1}}\cdots\hat{p}_{\alpha_{n}}$ (175)
where $G_{\alpha_{1}\cdots\alpha_{n}}$ is a matrix-valued symmetric tensor. The external gauge field can be introduced by replacing each $\hat{p}_{\alpha}$ with $\hat{\pi}_{\alpha}=\hat{p}_{\alpha}+e\mathcal{A}(\hat{x})$. Applying the Wigner transformation we get
$\hat{\pi}_{\alpha}\rightarrow$ $\,\pi_{\alpha}(x,p)=p_{\alpha}+e\mathcal{A}(x)$ (176)
$H_{0}(\hat{p})\rightarrow$ $\,{\sf H}_{0}(p)=\sum_{n=1}^{\infty}\frac{1}{n!}G_{\alpha_{1}\cdots\alpha_{n}}\,p_{\alpha_{1}}\cdots p_{\alpha_{n}}$ (177)
$H_{0}(\hat{\pi})\rightarrow$ $\,{\sf H}_{0\mathcal{A}}(x,p)=\sum_{n=1}^{\infty}\frac{1}{n!}G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}$ (178)
We don’t have any Moyal products in the second line because $p_{\alpha}\star p_{\beta}=p_{\alpha}p_{\beta}$ generally.
{mdframed}
Lemma
$G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}=$ $G_{\alpha_{1}\cdots\alpha_{n}}\left(\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n}}-\frac{e\hbar^{2}}{24}n(n-1)(n-2)\pi_{\alpha_{4}}\cdots\pi_{\alpha_{n}}\frac{\partial^{2}\mathcal{A}_{\alpha_{1}}(x)}{\partial x_{\alpha_{2}}\partial x_{\alpha_{3}}}\right)+\mathcal{O}(\hbar^{3})$
$=$ $\left(1-\frac{e\hbar^{2}}{24}\frac{\partial^{2}\mathcal{A}_{\mu}(x)}{\partial x_{\nu}\partial x_{\gamma}}\frac{\partial^{3}}{\partial\pi_{\mu}\partial\pi_{\nu}\partial\pi_{\gamma}}\right)G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n}}$ (179)
Proof: First, note $\left\{A\,\overset{\star}{,}\,B\right\}/2=AB-(\hbar^{2}/8)\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}A\partial^{2}_{\sigma\lambda}B+\mathcal{O}(\hbar^{3})$. Then, using the fact that $G_{\alpha_{1}\cdots\alpha_{n}}$ is symmetric under exchange of its indices we can get
$G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}=$ $\frac{1}{2}G_{\alpha_{1}\cdots\alpha_{n}}\,\left\{\pi_{\alpha_{2}}\star\cdots\star\pi_{\alpha_{n}}\,\overset{\star}{,}\,\pi_{\alpha_{1}}\right\}$
$=$ $G_{\alpha_{1}\cdots\alpha_{n}}\,\left((\pi_{\alpha_{2}}\star\cdots\star\pi_{\alpha_{n}})\pi_{\alpha_{1}}-\frac{\hbar^{2}}{8}\frac{\partial^{2}}{\partial p_{\nu}\partial p_{\gamma}}(\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n-1}})\frac{\partial^{2}\pi_{\alpha_{n}}}{\partial x_{\nu}\partial x_{\gamma}}\right)+\mathcal{O}(\hbar^{3})$
$=$ $G_{\alpha_{1}\cdots\alpha_{n}}\,\left((\pi_{\alpha_{2}}\star\cdots\star\pi_{\alpha_{n}})\pi_{\alpha_{1}}-(n-1)(n-2)\frac{e\hbar^{2}}{8}(\pi_{\alpha_{4}}\cdots\pi_{\alpha_{n}})\frac{\partial^{2}\mathcal{A}_{\alpha_{1}}}{\partial x_{\alpha_{2}}\partial x_{\alpha_{3}}}\right)+\mathcal{O}(\hbar^{3})$
Notice this sets up a recursive relation. The first term in the last line with the $\star$ products is the left-hand side up to differences in the symmetric tensor. If we continue this recursive relation we find
$G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}=$ $G_{\alpha_{1}\cdots\alpha_{n}}\,\left(\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n}}-\sum_{m=1}^{n-2}\frac{(n-m)!}{(n-m-2)!}\frac{e\hbar^{2}}{8}(\pi_{\alpha_{4}}\cdots\pi_{\alpha_{n}})\frac{\partial^{2}\mathcal{A}_{\alpha_{1}}}{\partial x_{\alpha_{2}}\partial x_{\alpha_{3}}}\right)+\mathcal{O}(\hbar^{3})$
Finally, we use the fact that $\sum_{m=1}^{n-2}\frac{(n-m)!}{(n-m-2)!}=\frac{1}{3}n(n-1)(n-2)$ and we get (179). Applying the lemma (179) we find
${\sf H}_{0\mathcal{A}}(x,p)={\sf H}_{0}(\pi)-\frac{e\hbar^{2}}{24}\frac{\partial^{3}{\sf H}_{0}(\pi)}{\partial\pi_{\mu}\partial\pi_{\nu}\partial\pi_{\gamma}}\frac{\partial^{2}\mathcal{A}_{\mu}(x)}{\partial x_{\nu}\partial x_{\gamma}}+\mathcal{O}(\hbar^{3}).$ (180)
We now consider when the Hamiltonian is not separable. We assume that the Hamiltonian is ordered such that the $\hat{x}$ and $\hat{p}$ operators are separately grouped together. In other words, we assume it takes the general form
$\hat{H}=\sum_{n}v_{n}(\hat{x})k_{n}(\hat{p})\Gamma_{n}+\rm{h.c.}$ (181)
where $\{v_{n}\},\{k_{n}\}$ are a set of polynomial functions and $\Gamma_{n}$ is a constant matrix. We now once again couple the system to an external gauge field by taking $\hat{p}\rightarrow\hat{\pi}$, and calculate the Wigner transformation. We know what the Wigner transformation of $k_{n}(\hat{\pi})$ must be from (180). Therefore,
${\sf H}_{\mathcal{A}}(x,p)=$ $\sum_{n}v_{n}(x)\star\left(k_{n}(\pi)-\frac{e\hbar^{2}}{24}\frac{\partial^{2}\mathcal{A}_{\mu}(x)}{\partial x_{\nu}\partial x_{\gamma}}\frac{\partial^{3}k_{n}(\pi)}{\partial\pi_{\mu}\partial\pi_{\nu}\partial\pi_{\gamma}}\right)\Gamma_{n}+\rm{h.c.}+\mathcal{O}(\hbar^{3})$
$=$ $\left(1-\frac{e\hbar^{2}}{24}\frac{\partial^{2}\mathcal{A}_{\mu}(x)}{\partial x_{\nu}\partial x_{\gamma}}\frac{\partial^{3}}{\partial\pi_{\mu}\partial\pi_{\nu}\partial\pi_{\gamma}}\right)\sum_{n}v_{n}(x)\star k_{n}(\pi)\Gamma_{n}+\rm{h.c.}+\mathcal{O}(\hbar^{3})$
$=$ $\left(1-\frac{e\hbar^{2}}{24}\frac{\partial^{2}\mathcal{A}_{\mu}(x)}{\partial x_{\nu}\partial x_{\gamma}}\frac{\partial^{3}}{\partial\pi_{\mu}\partial\pi_{\nu}\partial\pi_{\gamma}}\right){\sf H}(x,\pi)+\mathcal{O}(\hbar^{3})$ (182)
To get the second line, we simply move the derivatives outside of the sum since $k_{n}(\pi)$ is the only part that’s dependent on $\pi$. In addition, we recognize that $\sum_{n}v_{n}(x)\star k_{n}(\pi)\Gamma_{n}+\textrm{h.c.}=\sum_{n}v_{n}(x)\star k_{n}(p)\Gamma_{n}+\rm{h.c.}|_{p=\pi}={\sf H}(x,\pi)$. Therefore, (180) generalizes to arbitrary Hamiltonians (that can be expressed as a polynomial of $\hat{x}$ and $\hat{p}$.)
Appendix G Projection operator identities
The projection operator for the $n$-th band is defined as
$P_{n}\equiv U\star p_{n}\star U^{\dagger}$ (183)
where the matrix elements of $p_{n}$ are given by $(p_{n})_{ij}=\delta_{in}\delta_{jn}$. The projection operator is idempotent $P_{n}\star P_{n}=P_{n}$ and sum to identity $\sum_{n=1}^{N}P_{n}=I_{N}$.
The trace of the projection operator is not $1$. Rather,
$\mathrm{tr}[P_{n}]=1+\frac{\hbar}{2}\omega^{\alpha\beta}\Omega_{n,\alpha\beta}+\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\left(\Omega_{n,\alpha\sigma}\Omega_{n,\beta\lambda}-\frac{1}{2}\Omega_{n,\alpha\beta}\Omega_{n,\sigma\lambda}\right)+\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}g_{n,\beta\lambda}+\mathcal{O}(\hbar^{3})$ (184)
which can be derived using eq. 131.
Taking derivatives of the idempotence property $P_{n}\star P_{n}=P_{n}$ gives us
$\partial_{\alpha}P_{n}=\partial_{\alpha}P_{n}\star P_{n}+P_{n}\star\partial_{\alpha}P_{n},$ (185)
$\partial^{2}_{\alpha\sigma}P_{n}=\partial^{2}_{\alpha\sigma}P_{n}\star P_{n}+P_{n}\star\partial^{2}_{\alpha\sigma}P_{n}+\partial_{\alpha}P_{n}\star\partial_{\sigma}P_{n}+\partial_{\sigma}P_{n}\star\partial_{\alpha}P_{n},$ (186)
$\partial^{3}_{\alpha\sigma\mu}P_{n}=$ $\partial^{3}_{\alpha\sigma\mu}P_{n}\star P_{n}+P_{n}\star\partial^{3}_{\alpha\sigma\mu}P_{n}+\partial_{\alpha}P_{n}\star\partial^{2}_{\sigma\mu}P_{n}+\partial_{\sigma}P_{n}\star\partial^{2}_{\alpha\mu}P_{n}+\partial_{\mu}P_{n}\star\partial^{2}_{\alpha\sigma}P_{n}$ (187)
$+\partial^{2}_{\sigma\mu}P_{n}\star\partial_{\alpha}P_{n}+\partial^{2}_{\alpha\mu}P_{n}\star\partial_{\sigma}P_{n}+\partial^{2}_{\alpha\sigma}P_{n}\star\partial_{\mu}P_{n}.$
Using these three relations we can derive many useful identities. Below we list some useful trace identities.
$\mathrm{tr}[P_{n}\star\partial_{\alpha}P_{n}\star P_{n}]=0$ (188)
$\mathrm{tr}[\partial_{\alpha}P_{n}\partial_{\beta}P_{n}\partial_{\gamma}P_{n}]=0+\mathcal{O}(\hbar)$ (189)
$\mathrm{tr}[\partial_{\beta}P_{n}\star P_{n}]=i\hbar\omega^{\sigma\lambda}\partial_{\sigma}\mathcal{T}_{n,\lambda\beta}+\mathcal{O}(\hbar^{2})$ (190)
$\mathrm{tr}[\partial_{\alpha}P_{n}\partial_{\beta}P_{n}]=2g_{n,\alpha\beta}+\mathcal{O}(\hbar)$ (191)
$\mathrm{tr}[P_{n}\star\partial^{2}_{\beta\lambda}P_{n}\star P_{n}]=-2g_{n,\beta\lambda}$ (192)
$\mathrm{tr}[P_{n}\partial^{3}_{\beta\lambda\nu}P_{n}P_{n}]=-\partial_{\nu}g_{n,\beta\lambda}-\partial_{\lambda}g_{n,\nu\beta}-\partial_{\beta}g_{n,\nu\lambda}+\mathcal{O}(\hbar)$ (193)
$\mathrm{tr}[\partial_{\beta}P_{n}\partial^{2}_{\nu\lambda}P_{n}]=\partial_{\nu}g_{n,\beta\lambda}+\partial_{\lambda}g_{n,\nu\beta}-\partial_{\beta}g_{n,\nu\lambda}+\mathcal{O}(\hbar)$ (194)
Also look at Ref. [39] for a further discussion on projection-operator algebra.
Appendix H Projection operator representation of ${\sf H}$ and ${\sf F}$
In this section, we show that ${\sf H}$ and ${\sf F}$ can be expressed as
${\sf H}$ $=\sum_{n=1}^{N}P_{n}\star h_{n}\star P_{n}$ (195)
${\sf F}$ $=\sum_{n=1}^{N}P_{n}\star\underline{f}_{n}\star P_{n}+U\star\tilde{\mathcal{F}}\star U^{\dagger}$ (196)
where
$\underline{f}=\bigg{(}1-\frac{\hbar}{2}\omega^{\alpha\beta}\Omega_{\alpha\beta}-\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\bigg{(}\frac{1}{2}\Omega_{\alpha\sigma}\Omega_{\beta\lambda}-\frac{3}{4}\Omega_{\alpha\beta}\Omega_{\sigma\lambda}+\partial^{2}_{\alpha\sigma}g_{\beta\lambda}\bigg{)}\bigg{)}f-\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial_{\sigma}(\partial_{\alpha}fg_{\beta\lambda}).$ (197)
The Hamiltonian can be expressed as ${\sf H}=U\star\tilde{h}\star U^{\dagger}$. From this expression, it is straightforward to show that the Hamiltonian commutes with the projection operator, $\left[{\sf H}\,\overset{\star}{,}\,P_{n}\right]=0$. Since $\sum_{n=1}^{N}P_{n}=I_{n}$, we can write ${\sf H}=\sum_{n}P_{n}\star{\sf H}\star P_{n}=\sum_{n}U\star\tilde{h}_{n}p_{n}\star U^{\dagger}$. We now make an ansatz that there exists a $\underline{h}_{n}$ such that
$P_{n}\star\underline{h}_{n}\star P_{n}=U\star\tilde{h}_{n}p_{n}\star U^{\dagger}.$ (198)
We can take the trace of both sides and use eqs. 130 and 131 to solve for $\underline{h}$.
$\mathrm{tr}[P_{n}\star\underline{h}_{n}\star P_{n}]=\underline{h}_{n}\mathrm{tr}[P_{n}]+\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial_{\sigma}(\partial_{\alpha}\underline{h}_{n}g_{n,\beta\lambda})+\mathcal{O}(\hbar^{3})$ (199)
and
$\mathrm{tr}[U\star\tilde{h}_{n}p_{n}\star U^{\dagger}]=\tilde{h}_{n}+\hbar\omega^{\alpha\beta}\partial_{\alpha}(\tilde{h}_{n}A_{n\beta})+\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}(\tilde{h}_{n}\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\}_{nn})+\mathcal{O}(\hbar^{3})$ (200)
If we then solve for $\underline{h}$ we find that $\underline{h}=h$ which gives us eq. 195.
Unlike the Hamiltonian, the distribution function is expressed as the sum of two parts ${\sf F}=U\star\tilde{f}\star U^{\dagger}+U\star\tilde{\mathcal{F}}\star U^{\dagger}$. The first term can be expressed in the same way as the Hamiltonian since $\tilde{f}$ is diagonal. Therefore, we define ${\sf F}=\sum_{n=1}^{N}P_{n}\star\underline{f}_{n}\star P_{n}+U\star\tilde{\mathcal{F}}\star U^{\dagger}$. We can recycle the fact that $\underline{h}=h$ to determine that
$\underline{f}=\tilde{f}_{n}+\hbar\omega^{\alpha\beta}\partial_{\alpha}\tilde{f}_{n}(A_{n\beta}+\hbar\omega^{\sigma\lambda}A_{n\lambda}\partial_{\sigma}A_{n\beta})-\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{f}_{n}\left(\frac{1}{4}\left\{\Lambda_{\beta}\,,\Lambda_{\lambda}\right\}_{nn}-A_{n\beta}A_{n\lambda}\right)+\mathcal{O}(\hbar^{3})$ (201)
by simply taking the definition for $h$ and replacing $\tilde{h}$ with $\tilde{f}$. Next, we use eq. 136 to express $\tilde{f}$ using $f$ and obtain eq. 197.
Appendix I Expressing electric current using $h,f$ (eq. 47)
Here, we outline the derivation of eq. 47. First, we generalize the current given by eq. 46 to a phase-space current. This is straightforward and we define
$\mathcal{J}_{\mu}(x,p)\equiv\omega^{\mu\nu}\operatorname{Re}\mathrm{tr}[\partial_{\nu}{\sf H}\star{\sf F}]+\frac{\hbar^{2}}{12}\omega^{\mu\nu}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\mathrm{tr}[\partial^{3}_{\nu\beta\lambda}{\sf H}{\sf F}]+\mathcal{O}(\hbar^{3})$ (202)
such that $J_{i}(x)=-e\int_{p}\mathcal{J}_{x_{i}}(x,p)$. Assuming the off-diagonal components of the distribution function, $\tilde{\mathcal{F}}$, are negligible, we can write ${\sf F}=\sum_{n}P_{n}\star\underline{f}_{n}\star P_{n}$ which is the projection operator representation discussed in appendix H. Substituting this into the expression for the phase-space current and using eq. 130, we find
$\mathcal{J}_{\mu}=$ $\omega^{\mu\nu}\sum_{n}\underline{f}_{n}V_{n,\nu}+\omega^{\mu\nu}\hbar\omega^{\alpha\beta}\partial_{\alpha}\sum_{n}\underline{f}_{n}V_{n,\nu\beta}+\frac{\hbar^{2}}{12}\omega^{\mu\nu}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\sum_{n}\underline{f}_{n}V_{n,\nu\beta\lambda}+\mathcal{O}(\hbar^{3})$ (203)
where
$V_{n,\nu}\equiv\mathrm{tr}[P_{n}\star\partial_{\nu}{\sf H}\star P_{n}]$ (204)
$V_{n,\nu\beta}\equiv-\operatorname{Im}\mathrm{tr}[\partial_{\beta}P_{n}\star\partial_{\nu}{\sf H}\star P_{n}]$ (205)
$V_{n,\nu\beta\lambda}\equiv-\frac{1}{2}\partial^{3}_{\nu\beta\lambda}h_{n}+\frac{1}{2}\partial_{\lambda}\mathrm{tr}[\partial_{\beta}P_{n}\partial_{\nu}{\sf H}]+\frac{1}{2}\partial_{\beta}\mathrm{tr}[\partial_{\lambda}P_{n}\partial_{\nu}{\sf H}]-2\mathrm{tr}[\partial^{2}_{\beta\lambda}P_{n}\partial_{\nu}{\sf H}]+6\operatorname{Re}\mathrm{tr}[\partial_{\beta}P_{n}\partial_{\nu}{\sf H}\partial_{\lambda}P_{n}]$ (206)
We can now rewrite each of these traces by substituting ${\sf H}=\sum_{n}P_{n}\star h_{n}\star P_{n}$ and using the trace identities of the projection operator listed in appendix G. We find
$\begin{split}V_{\nu}=&\partial_{\nu}h\left(1+\frac{\hbar}{2}\omega^{\alpha\beta}\Omega_{\alpha\beta}+\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\left(\Omega_{\alpha\sigma}\Omega_{\beta\lambda}-\frac{1}{2}\Omega_{\alpha\beta}\Omega_{\sigma\lambda}\right)+\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}g_{\beta\lambda}\right)\\ &-\hbar\omega^{\alpha\beta}\partial_{\alpha}h\Omega_{\nu\beta}+\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\left(\frac{1}{2}\partial^{2}_{\alpha\sigma}h\partial_{\nu}g_{\beta\lambda}+\partial^{3}_{\nu\alpha\sigma}hg_{\beta\lambda}+\partial^{2}_{\nu\alpha}h\partial_{\sigma}g_{\beta\lambda}\right)+\mathcal{O}(\hbar^{3})\end{split}$ (207)
$V_{\nu\beta}=-m_{\nu\beta}-\hbar\omega^{\sigma\lambda}(\partial^{2}_{\nu\sigma}hg_{\lambda\beta}+\partial_{\nu}h\partial_{\sigma}g_{\lambda\beta})+\frac{\hbar}{2}\omega^{\sigma\lambda}\partial_{\sigma}h(\partial_{\nu}g_{\beta\lambda}-\partial_{\beta}g_{\lambda\nu}+\partial_{\lambda}g_{\nu\beta})+\mathcal{O}(\hbar^{2})$ (208)
$V_{\nu\beta\lambda}=-\frac{1}{2}\partial^{3}_{\nu\beta\lambda}h+12\partial_{\nu}hg_{\beta\lambda}+2\partial_{\nu}c_{\beta\lambda}-\partial_{\lambda}c_{\beta\nu}-\partial_{\beta}c_{\lambda\nu}+4c_{\beta\lambda;\nu}-2c_{\beta\nu;\lambda}-2c_{\lambda\nu;\beta}+\mathcal{O}(\hbar)$ (209)
where we defined $m$ and $c$’s to be diagonal matrices whose $n$-th diagonal elements are given by
$m_{n,\nu\beta}\equiv\operatorname{Im}\mathrm{tr}[\partial_{\nu}P_{n}\star({\sf H}-h_{n}I_{N})\star\partial_{\beta}P_{n}]$ (210)
$c_{n,\nu\beta}\equiv\operatorname{Re}\mathrm{tr}[\partial_{\nu}P_{n}\star({\sf H}-h_{n}I_{N})\star\partial_{\beta}P_{n}]$ (211)
$c_{n,\beta\lambda;\nu}\equiv\operatorname{Re}\mathrm{tr}[\partial^{2}_{\beta\lambda}P_{n}\star({\sf H}-h_{n}I_{N})\star\partial_{\nu}P_{n}]$ (212)
Next, we express $\underline{f}$ using $f$ by substituting
$\underline{f}=\bigg{(}1-\frac{\hbar}{2}\omega^{\alpha\beta}\Omega_{\alpha\beta}-\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\bigg{(}\frac{1}{2}\Omega_{\alpha\sigma}\Omega_{\beta\lambda}\\ -\frac{3}{4}\Omega_{\alpha\beta}\Omega_{\sigma\lambda}+\partial^{2}_{\alpha\sigma}g_{\beta\lambda}\bigg{)}\bigg{)}f-\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial_{\sigma}(\partial_{\alpha}fg_{\beta\lambda}).$ (213)
We then find
$\begin{split}\mathcal{J}_{\mu}=&\omega^{\mu\nu}\mathrm{tr}\left[f\left(\partial_{\nu}h-\hbar\omega^{\alpha\beta}\partial_{\alpha}h\Omega_{\nu\beta}(1-\frac{\hbar}{2}\omega^{\sigma\lambda}\Omega_{\sigma\lambda})+\frac{\hbar^{2}}{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}h\partial_{\nu}g_{\beta\lambda}\right)\right]\\ &+\hbar\omega^{\mu\nu}\omega^{\alpha\beta}\partial_{\alpha}\mathrm{tr}[f(b^{s}_{\nu\beta}+b^{a}_{\nu\beta})]+\frac{\hbar^{2}}{12}\omega^{\mu\nu}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\mathrm{tr}[f(b^{s}_{\nu\beta\lambda}+b^{m}_{\nu\beta;\lambda}+b^{m}_{\nu\lambda;\beta})]\end{split}$ (214)
where
$b_{\nu\beta}^{s}\equiv\frac{\hbar}{2}\omega^{\sigma\lambda}(\partial_{\sigma}h\partial_{\lambda}g_{\nu\beta}+\partial^{2}_{\sigma\nu}hg_{\beta\lambda}+\partial^{2}_{\sigma\beta}hg_{\nu\lambda})$ (215)
$b^{a}_{\nu\beta}\equiv-m_{\nu\beta}\left(1-\frac{\hbar}{2}\omega^{\sigma\lambda}\Omega_{\sigma\lambda}\right)-\frac{\hbar}{2}\omega^{\sigma\lambda}\left(\partial_{\beta}(\partial_{\sigma}hg_{\nu\lambda})-\partial_{\nu}(\partial_{\sigma}hg_{\beta\lambda})\right)$ (216)
$b_{\nu\beta\lambda}^{s}\equiv-\frac{1}{2}\partial^{3}_{\nu\beta\lambda}h$ (217)
$b^{m}_{\nu\lambda;\beta}\equiv\partial_{\nu}c_{\beta\lambda}-\partial_{\lambda}c_{\beta\nu}+2c_{\beta\lambda;\nu}-2c_{\beta\nu;\lambda}$ (218)
Then, by calculating $J_{i}\equiv-e\int_{p}\mathcal{J}_{x_{i}}$ we get eq. 47.
Appendix J Chiral anomaly effect
The equilibrium real-space current has been shown in (54) to be divergence-free in the absence of external fields. Here we want to demonstrate that it will break down in the presence of external fields at a Weyl node, known as the chiral anomaly effect. In this section, we focus on the $O(\hbar)$ contributions.
According to Section 2.9, the additional contribution to the equilibrium real-space current in the presence of a uniform magnetic field is due to $\omega_{B}^{p_{i}p_{j}}\neq 0$, where we relabeled the kinetic momentum $\pi_{i}\rightarrow p_{i}$. Such a constribution is given by
$J^{\prime}_{i,\mathrm{eq}}=-e\hbar\omega_{B}^{p_{j}p_{k}}\int_{p}\mathrm{tr}\left[\left(-\partial_{p_{j}}h\Omega_{p_{i}p_{k}}+\frac{1}{2}\Omega_{p_{j}p_{k}}\partial_{p_{i}}h\right)n_{\mathrm{F}}\right],$ (219)
whose divergence is given by
$\partial_{x_{i}}J^{\prime}_{i,\mathrm{eq}}$ $=-e\hbar\omega^{p_{j}p_{k}}_{B}\int_{p}\mathrm{tr}\left[\left(-\partial_{p_{j}}h\partial_{x_{i}}\Omega_{p_{i}p_{k}}+\partial_{x_{i}}h\partial_{p_{j}}\Omega_{p_{i}p_{k}}+\frac{1}{2}\partial_{x_{i}}\Omega_{p_{j}p_{k}}\partial_{p_{i}}h-\frac{1}{2}\partial_{p_{i}}\Omega_{p_{j}p_{k}}\partial_{x_{i}}h\right)n_{\mathrm{F}}\right],$ (220)
where we used integration by part. In $d=2$, we can write $\omega_{B}^{p_{j}p_{k}}=-eB\epsilon^{jk}$ and $\Omega_{p_{i}p_{j}}=\Omega\epsilon_{ij}$, and we find
$\partial_{x_{i}}J^{\prime}_{i,\mathrm{eq}}$ $=0.$ (221)
In $d=3$, we can write $\omega_{B}^{p_{j}p_{k}}=-eB_{l}\epsilon^{jkl}$ and $\Omega_{p_{i}p_{j}}=\Omega_{p}^{k}\epsilon_{ijk}$, and we find
$\partial_{x_{i}}J^{\prime}_{i,\mathrm{eq}}$ $=e^{2}\hbar B_{l}\int_{p}(\partial_{p_{k}}h\partial_{x_{l}}\Omega_{p}^{k}-\partial_{x_{l}}h\partial_{p_{k}}\Omega_{p}^{k})n_{\mathrm{F}}.$ (222)
For a regular (twice differentiable) Hamiltonian, $\Omega_{p}^{i}$ will also be regular in real space, thus the only non-trivial term would be the second term; if $\Omega_{p}^{i}$ is regular in momentum space, however, $\nabla_{p}\cdot\Omega_{p}=\nabla_{p}\cdot(\nabla_{p}\times A_{p})=0$. At a Weyl node, $\Omega_{p}^{i}=\pm p_{i}/2p^{3}$, thus, under an external electric field $\partial_{x_{i}}h=-eE_{i}$, we find the breakdown of charge conservation
$\partial_{x_{i}}J^{\prime}_{i,\mathrm{eq}}=\pm\frac{e^{3}\hbar}{4\pi^{2}}(E\cdot B),$ (223)
which is because at the gapless Weyl node, there is a Berry monopole $\nabla_{p}\cdot\Omega_{p}=2\pi\delta^{(3)}(p)$.
Appendix K Deriving $f_{r}$ for the relaxation time approximation (eq. 117)
Here, we derive $f_{r}$ given by eq. 117, the distribution function which $f$ relaxes to under the relaxation time approximation when the Hamiltonian takes the form ${\sf H}={\sf H}_{0}(p)-eV(x)I_{N}$. The relaxation time approximation of the collision integral for the original Liouville-von Neumann equation given by eq. 11 is $-\tau^{-1}({\sf F}-{\sf F}_{r})$ where ${\sf F}_{r}={\sf F}_{\textrm{eq}}=U^{(0)}n_{F}(\varepsilon)U^{(0)\dagger}$. To calculate $f_{r}$, we need to first find an explicit expression for $\tilde{f}_{r}=U^{\dagger}\star{\sf F}_{r}\star U$. To do this, we use $T=\hbar U^{\dagger}\star\partial_{\hbar}U$ which we defined when studying flow equations in appendix D. Recall, $T=Y+\frac{\hbar}{4}\omega^{\alpha\beta}\partial_{\alpha}\Lambda_{\beta}$ where $Y$ is the anti-Hermitian part and the second term is the Hermitian part of $T$. If we expand $T$ in powers of $\hbar$ we find
$T=$ $(U^{(0)\dagger}+U^{(1)\dagger}+\cdots)\star(U^{(1)}+2U^{(2)}+\cdots)$ (224)
$=$ $U^{(0)\dagger}U^{(1)}+2U^{(0)\dagger}U^{(2)}+U^{(1)^{\dagger}}U^{(1)}+\frac{i\hbar}{2}\omega^{\alpha\beta}\partial_{\alpha}U^{(0)\dagger}\partial_{\beta}U^{(1)}+\mathcal{O}(\hbar^{3})$ (225)
For the Hamiltonian we are considering, $T^{(1)}=Y^{(1)}$ since $\Lambda_{x_{i}}=0$ and $\Lambda_{p_{i}}^{(0)}=\Lambda_{p_{i}}^{(0)}(p)$. Therefore, $U^{(0)\dagger}U^{(1)}=Y^{(1)}$. In addition,
$U^{(0)\dagger}U^{(2)}=\frac{1}{2}\left(T^{(2)}+(Y^{(1)})^{2}+\frac{\hbar}{2}\Lambda_{p_{i}}^{(0)}\partial_{x_{i}}Y^{(1)}\right)$ (226)
so the Hermitian part of $U^{(0)\dagger}U^{(2)}$ is
$\frac{1}{2}(U^{(0)\dagger}U^{(2)}+U^{(2)\dagger}U^{(0)})=-\frac{\hbar}{2}\left[\partial_{x_{i}}Y^{(1)}\,,\Lambda_{p_{i}}^{(0)}\right]+(Y^{(1)})^{2},$ (227)
where we used the fact that $A^{(1)}_{p_{i}}=\left[\Lambda_{p_{i}}^{(0)}\,,Y^{(1)}\right]^{d}$.
Let’s now expand $\tilde{f}_{r}$ to second order in $\hbar$. At zeroth order, we trivially find $\tilde{f}_{r}^{(0)}=n_{F}(\varepsilon)$. To first order it is given by $\tilde{f}_{r}^{(1)}=\operatorname{diag}(T^{(1)}+T^{(1)\dagger})n_{F}(\varepsilon)=0$. The second-order contribution is
$\tilde{f}_{r}^{(2)}=$ $U^{(2)\dagger}{\sf F}_{r}U^{(0)}+U^{(0)\dagger}{\sf F}_{r}U^{(2)}+U^{(1)\dagger}{\sf F}_{r}U^{(1)}$ (228)
$+\frac{i\hbar}{2}(\partial_{x_{i}}U^{(1)\dagger}\partial_{p_{i}}{\sf F}_{r}U^{(0)}-U^{(0)\dagger}\partial_{p_{i}}{\sf F}_{r}\partial_{x_{i}}U^{(1)}+\partial_{x_{i}}U^{(1)\dagger}{\sf F}_{r}\partial_{p_{i}}U^{(0)}-\partial_{p_{i}}U^{(0)\dagger}{\sf F}_{r}\partial_{x_{i}}U^{(1)})$
$=$ $(U^{(2)\dagger}U^{(0)}+U^{(0)\dagger}U^{(2)})^{d}n_{F}(\varepsilon)-(Y^{(1)}n_{F}(\varepsilon)Y^{(1)})^{d}+\frac{\hbar}{2}\left[\partial_{x_{i}}Y^{(1)}\,,\Lambda_{p_{i}}^{(0)}\right]^{d}n_{F}(\varepsilon)$ (229)
$=$ $\left(Y^{(1)}\left[Y^{(1)}\,,n_{F}(\varepsilon)\right]\right)^{d}$ (230)
Here, we used $(\cdot)^{d}$ as short hand for $\operatorname{diag}(\cdot)$. We took the diagonal of the right-hand side since $\tilde{f}_{r}$ is assumed to be diagonal. Since $(Y^{(1)})_{nm}=-\frac{e}{\varepsilon_{n}-\varepsilon_{m}}(\Lambda^{(0)}_{p_{i}})_{nm}\partial_{x_{i}}V$ for $n\neq m$, if we take just the $n$-th element of $\tilde{f}_{r}^{(2)}$, we see we can rewrite it as
$\tilde{f}_{r,n}^{(2)}=-e^{2}\sum_{m\neq n}(\Lambda_{p_{i}}^{(0)})_{nm}(\Lambda_{p_{j}}^{(0)})_{mn}\frac{n_{F,n}-n_{F,m}}{(\varepsilon_{n}-\varepsilon_{m})^{2}}\partial_{x_{i}}V\partial_{x_{j}}V.$ (231)
Now, $f_{r}$ is $f_{r}=\tilde{f}_{r}+\hbar\tilde{f}_{r}\partial_{x_{i}}A_{p_{i}}+\mathcal{O}(\hbar^{3})$ from eq. 26. Therefore,
$f_{r,n}=n_{F,n}(1-e\hbar^{2}T_{n,ij}\partial^{2}_{x_{i}x_{j}}V)-e^{2}\sum_{m\neq n}\langle\partial_{p_{i}}u_{n}|u_{m}\rangle\langle u_{m}|\partial_{p_{j}}u_{n}\rangle\frac{n_{F,n}-n_{F,m}}{(\varepsilon_{n}-\varepsilon_{m})^{2}}\partial_{x_{i}}V\partial_{x_{j}}V$ (232)
where we used $(\Lambda^{(0)}_{p_{i}})_{nm}=-i\langle u_{n}|\partial_{p_{i}}u_{m}\rangle$.
Appendix L Density-density correlation function
The density-density correlation function for non-interacting multi-band fermions is given by
$\widehat{C}^{(2)}(q,\omega)=$ $i\hbar\sum_{\begin{subarray}{c}n,m=1\\ n\neq m\end{subarray}}^{N}\int_{k}(U^{(0)\dagger}(k-q)U^{(0)}(k))_{nm}(U^{(0)\dagger}(k)U^{(0)}(k-q))_{mn}\frac{n_{F,n}(k-q)-n_{F,m}(k)}{\hbar\omega-\varepsilon_{m}(k)+\varepsilon_{n}(k-q)}$ (233)
$\approx$ $-i\sum_{n=1}^{N}\int_{k}\frac{\hbar^{2}n_{F,n}^{\prime}\vec{v}\cdot\vec{q}\,(1-{\sf g}_{n,ij}q_{i}q_{j})}{\hbar\omega-\varepsilon_{n}(k)+\varepsilon_{n}(k-q)}+i\hbar\sum_{\begin{subarray}{c}n,m=1\\ n\neq m\end{subarray}}^{N}\int_{k}(n_{F,n}-n_{F,m})\frac{({\sf\Lambda}_{i})_{nm}({\sf\Lambda}_{j})_{mn}}{\hbar\omega-\varepsilon_{m}+\varepsilon_{n}}q_{i}q_{j}+\mathcal{O}(q^{3})$
where $n_{F,n}=n_{F}(\varepsilon_{n})$. The last line was obtained by expanding to second-order in $q$. In the low-frequency limit $\omega\ll\omega_{\textrm{gap}}$, the frequency in the denominator of the last term goes to zero and this term can be expressed using ${\sf t}_{ij}$. A straightforward calculation shows that this expression reduces to (127a). However, the full expression (233) is capable of describing an insulator, for which (128) vanishes. In particular, the interband static structure factor due to the last term in (233) is given by
$S_{q}^{\mathrm{inter}}=q_{i}q_{j}\int_{k}{\sf g}_{ij}^{\mathrm{occ}}(k),$ (234)
where ${\sf g}_{ij}^{\mathrm{occ}}\equiv\sum_{m>n,n\in\mathrm{occ}}({\sf\Lambda}_{i})_{nm}({\sf\Lambda}_{j})_{mn}$ is the quantum metric of the occupied bands [35].